### v1

# How Good is Your Wikipedia?

Abstract
Wikipedia’s perceived high quality and broad language coverage have established it as a fundamental resource in multilingual NLP. In the context of low-resource languages, however, these quality assumptions are increasingly being scrutinised. This paper critically examines the data quality of Wikipedia in a non-English setting by subjecting it to various quality filtering techniques, revealing widespread issues such as a high percentage of one-line articles and duplicate articles. We evaluate the downstream impact of quality filtering on Wikipedia and find that data quality pruning is an effective means for resource-efficient training without hurting performance, especially for low-resource languages. Moreover, we advocate for a shift in perspective from seeking a general definition of data quality towards a more language- and task-specific one. Ultimately, we aim for this study to serve as a guide to using Wikipedia for pretraining in a multilingual setting. We also introduce a package that implements our data cleaning, tokenization, and model training pipeline. [^1] https://github.com/akulmizev/texieve
**footnotetext: Equal contribution
1 Introduction
Wikipedia is an invaluable resource in the field of NLP. Much of its usefulness stems from its open-source nature, through which a community of authors, editors, and administrators can collaborate on the documentation of an ever-changing world. Such a framework ensures that information contained therein is not only current, but also well-written, accurate, and relevant to any given subject. For these reasons, Wikipedia has been employed for a myriad of NLP applications, ranging from language model training and benchmarking (Merity et al., 2016), question-answering (Hewlett et al., 2016; Rajpurkar et al., 2016; Joshi et al., 2017) and knowledge-base creation (Vrandečić and Krötzsch, 2014; Lehmann et al., 2024), among many others. Wikipedia is a staple resource in multilingual NLP, where it is commonly employed for pretraining multilingual language models (LMs) (Devlin et al., 2019; Conneau and Lample, 2019) and creating benchmarks for downstream tasks, such as named-entity recognition Rahimi et al. (2019) and machine translation Schwenk et al. (2021).
An increasing awareness of the importance of data quality has led to data filtering practices finding a more permanent place in the NLP pipeline, whereby Wikipedia’s encyclopaedic domain and culture of rigorous community moderation mean that it is typically assumed to be a ’high quality’ reference in this process (e.g., Wenzek et al., 2020). However, this assumption, while generally reliable for English and select higher-resourced languages, of Wikipedia as ‘high quality’, is now being challenged in the multilingual setting Kreutzer et al. (2022a). Wikipedias for lower-resourced languages can contain unnatural text written by non-native speakers, machine generated text, and other noise, which can be incomprehensible to native speakers Alabi et al. (2020); Lent et al. (2024). Poor data quality for low-resource languages can subsequently lead to poorly performing and untrustworthy models, ultimately resulting in unusable technologies for speakers Held et al. (2023); Nicholas and Bhatia (2023); Durmus et al. (2024).
In this study, we critically examine the notion of Wikipedia as ‘high quality’ in a non-English setting. To this end, we compare common techniques for data filtering and how they perform on Wikipedia, particularly with attention to lower-resourced settings. We demonstrate the lack of apparent data quality in Wikipedia by showing that a lot of non-English Wikipedias contain foreign characters and duplicate articles, which we remove in a primary filtering step, and further apply corpus-driven heuristic quality filters. Downstream evaluation shows that these quality filters are not detrimental to performance, and thus can be used as a means for more resource-efficient training. Ultimately, we argue for a shift towards task and language specific approaches to data quality rather than seeking a general definition, especially when working with lower-resource languages. Our contributions are as follows:
  1. 1.
    We examine the effects of standard data cleaning techniques across non-English Wikipedias.
  2. 2.
    We evaluate the effect of data quality filtering on downstream performance, particularly in a lower-resource setting.
  3. 3.
    We demonstrate the importance of language and task specific approaches to data quality filtering.
With this, we add to the field’s larger debate about data quantity versus data quality. We aim for this study to be a starting guide to using Wikipedia as a resource for pretraining in a multilingual setting.
2 Related Work
2.1 Data Filtering
Recent works in large language modelling have demonstrated the benefits of data filtering instead of pretraining over unpruned data dumps Wenzek et al. (2020); Raffel et al. (2023); Penedo et al. (2024); 01.AI et al. (2024). In this context, data filtering typically involves a combination of techniques, [^2] For a comprehensive survey of strategies for data filtering for LLMs, we refer readers to Albalak et al. (2024). ranging from simple heuristics and document deduplication to learned filters that estimate quality. Examples of heuristics from C4 include removing document containing code (e.g., Javascript), boiler plate legal policies (e.g., ‘Terms and Conditions’), or inappropriate terms, amongst other hand-crafted rules Raffel et al. (2023). Meanwhile, a popular technique for deduplication is MinHash, as it is easily scalable across CPU nodes Penedo et al. (2024); others combine MinHash with downstream clustering-based deduplication to filter more aggressively Tirumala et al. (2023). In addition to heuristics and deduplication, 01.AI et al. (2024) leverage learned filters (i.e., a quality classifier and cluster-based filters), which are designed to preserve documents resembling a ‘high quality’ reference corpus – and typically Wikipedia Wenzek et al. (2020); Together Computer (2023).
2.2 Wikipedia in Multilingual NLP
Wikipedia is the backbone of many massively multilingual benchmark datasets, such as WikiAnn Pan et al. (2017), XQuAD Artetxe et al. (2020), MLQA Lewis et al. (2020), TyDiQA-GoldP Clark et al. (2020), and BUCC Zweigenbaum et al. (2017), as well as a common source of pretraining data for popular multilingual LMs like mBERT Devlin et al. (2019) and XLM Lample and Conneau (2019). Even the newer and larger LMs (e.g., LLaMA Touvron et al. (2023a)), pretrained over enormous web corpora, can be critically dependent on Wikipedia. For example, CCNet’s language identification classifier was trained over the online encyclopedia Wenzek et al. (2020). Indeed, Wikipedia can be found in every part of the NLP pipeline from language identification to pretraining and fine-tuning, and conclusions drawn from massively multilingual studies are thus largely inextricable from Wikipedia Hu et al. (2020).
As a consequence of its ubiquity in NLP, Wikipedia is attracting increased scrutiny in terms of its data quality. A large scale quality audit of large web crawled datasets finds WikiMatrix Schwenk et al. (2021) has a number of deprecated, incorrect, or otherwise mislabelled language codes, among the 86 languages included Kreutzer et al. (2022a). Their additional manual audit of samples uncovers that quality gradually declines in correlation with resourcedness. Lignos et al. (2022) identify similar shortcomings for lower-resourced languages in WikiAnn and even Wikidata. [^3] https://wikidata.org Studies on selected low-resource languages further confirm this finding. For example, Alabi et al. (2020) observe that the Yorùbá and Twi Wikipedias are ostensibly written by non-native speakers and lack the necessary diacritics and variation to be relevant to speakers. Meanwhile, Lent et al. (2024) find that sentences from Wikipedias for Creole languages are often incomprehensible by speakers and require manual correction before being used to create benchmarks.
2.3 Data Quality versus Data Quantity
The notion of data quality is inherently subjective. Quantitatively, quality is typically measured by the relative boosting performance on a downstream task. Qualitatively, quality can be understood to mean samples which are accepted as valid by proficient speakers of a language, or samples that do not promote toxic or offensive sentiments Kreutzer et al. (2022b); van Noord et al. (2024). These definitions can sometimes be at odds with one another. For example, Longpre et al. (2024) show that a smaller, carefully filtered dataset can improve performance on downstream tasks in contrast to a larger, unfiltered dataset, but that the resulting model becomes more likely to generate toxic content. There are also diminishing gains from training on ever increasing sizes of data Warstadt et al. (2023). For pretraining LLMs, C4 (745GB) marginally outperforms Wikipedia (16GB) over benchmarks like GLUE, SGLEU, CNNDM, and SQuAD, as many of these benchmarks were sourced from Wikipedia, suggesting that quantity is less important than domain Raffel et al. (2023). For task-specific fine-tuning of large LMs, smaller, manually annotated datasets also lead to better performance than massive, open-source datasets Touvron et al. (2023b); Zhou et al. (2023); 01.AI et al. (2024); Gemini Team et al. (2024); Üstün et al. (2024). That said, data filtering is not a universally accepted practice; critics of filtering call attention to the risks of manipulating data, rather than using it ‘as it is’ Goldberg (2021). We refer readers to Rogers (2021) for a focused conversation on the benefits and drawbacks of data-curation within the context of training NLP models, both for general pre-training and for downstream tasks.
3 Is Wikipedia ‘High Quality’?
Figure 1: Proportion of articles and characters filtered from non-English Wikipedias with primary filtering.
Albalak et al. (2024) outline a set of best practices for quality data selection and filtering. In this work, we examine the effects of their proposed model-agnostic filtering strategies [^4] We exclude classifier-based and perplexity-based filtering strategies measuring quality relative to a selection of data., as applied to Wikipedia. Conceptually, we can recognize such filters as belonging to two distinct categories based on their underlying objective: 1) procedures for removing undesirable material from Wikipedia, such as foreign scripts and duplicate articles; and 2) heuristics that aim to enhance the existing data distribution by removing otherwise lower quality text according to a given metric (e.g. article length in characters). For the first, we posit that if this step removes a substantial amount of text from a given Wikipedia dump, the removed content was not of high quality to begin with. We refer to this step as primary filtering. For the second, we aim to further refine Wikipedia articles via heuristic filters for even higher quality.
3.1 Primary Filtering
Language Filtering
Applying an existing language identification model to Wikipedia is a challenge due to data contamination – most models supporting all of Wikipedia’s languages have generally also been trained on Wikipedia (see e.g. Kargaran et al., 2023). Thus, we employ script-based filtering as a proxy for language identification by filtering characters in the Wikipedia that are not in the officially documented script(s) [^5] For example, the Konkani Wikipedia has been documented to be written in 3 scripts – Devanagari, Latin and Kannada. We filter all characters that do not belong to any of these scripts. for the concerned language. [^6] https://en.wikipedia.org/wiki/List_of_Wikipedias Figure 1 demonstrates that 3.37% of characters in non-English Wikipedias are affected by this. Some Wikipedias even contain entire articles written in another language, generally English. [^7] We acknowledge that this methodology would not be effective for such English articles in non-English Wikipedias with Latin script. However, this would require a language ID model that has not been trained on Wikipedia (and therefore may have been trained on these articles as the ground truth.) For example, 6% of Assamese Wikipedia consists of foreign characters, where numerous articles are written almost entirely in English Wilson (2019). The implications of this are inevitably made clear when training a tokeniser on this Wikipedia, where much of the vocabulary budget can be used up by foreign characters.
Exact-match Deduplication
We employ hashing to identify and remove articles which have the exact same text, but a different unique ID or title. We find that some Wikipedias are strongly affected by this process. For example, the Yorùbá Wikipedia loses almost 60% of its articles. Typically, articles removed by exact-match deduplication are placeholders containing filler or templatic text and generally contain little usable information. Figure 1 shows that 1.01% of non-English Wikipedias articles are exact duplicates.
Minhash Deduplication
While exact-match is able to filter identical documents, it cannot account for partial duplicates — i.e., sets of documents containing mostly identical text, but varying in terms of single sentences, paragraphs, or even hyperlinks. For example, the Yorùbá articles for the Baluch people and the Baluchi language are largely identical, with the latter containing an additional language data panel. To capture such cases, we perform minhash deduplication with locality sensitive hashing (LSH) and discount any article that returns a Jaccard similarity of 0.85. We find that a staggering 28.33% of all non-English Wikipedia articles (and 8.18% characters) are affected by this procedure — largely consisting of Wikipedias that have been known to be majority bot generated, such as Cebuano and Swedish. [^8] https://blog.datawrapper.de/wikipedia-articles-written-by-a-bot/
3.2 Heuristic Filtering
While the aforementioned primary filtering steps can reliably discount many undesirable documents, additional constraints may be considered in order to further refine the data selection process. To this end, we adopt what Albalak et al. (2024) categorise as heuristic filtering: a process in which document-level statistics are calculated in order to remove low quality data beyond a given threshold. Heuristics, in this sense, refer to simple metrics that aim to numerically characterise document content — for example, the number of characters contained in a given document. The red-pajama-data-v2 (Together Computer, 2023) dataset cleaning pipeline offers a collection of many such heuristics, culled from previous data curation efforts (Rae et al., 2021; Raffel et al., 2020; Longpre et al., 2024; Penedo et al., 2023). In theory, if Wikipedia is assumed to be a ‘high quality’ data source, it should be robust to these heuristics and not suffer from large-scale document removal.
To determine if this is the case, we explore natural language heuristics that are relevant to the encyclopedic nature of Wikipedia and easily applicable in a multilingual setting. [^9] For example, rps_doc_stop_word_fraction is not employed because it is difficult to principally obtain stopword lists for all the languages in Wikipedia. These are rps_doc_frac_unique_words (type-token ratio), rps_doc_unigram_entropy, and rps_doc_word_count — as collected by RedPajama. We also experiment with several adaptations to these metrics, such as utilising trigrams or characters as the basic unit of measurement (instead of whitespace-separated words). Naturally, some of these metrics demonstrate a very high correlation with each other; for example, absolute counts of characters, words, and trigrams typically exhibit a very high positive pairwise correlation across documents ($\rho>0.95$). As such, we decide to group all metrics into three classes: absolute, relative, and entropy. We list all metrics and their associated classes in Table 1.
Heuristic-based Quality Filters
Class Name Calculate…
Absolute length length in characters.
unique_trigrams total no. of unique trigrams.
unique_words total no. of unique words.
Relative frac_unique_trigrams unique trigrams / total trigrams
frac_unique_words unique words / total words
Entropy trigram_entropy $-\sum_{t\in T}\left(\frac{\text{count}(t)}{N}\right)\cdot\log_{2}\left(\frac{\text{count}(t)}{N}\right)$
unigram_entropy
Table 1: The quality heuristics applied to Wikipedia, calculated for each article. For entropy, $t$ is a unique n-gram in a set of $T$ unique document n-grams, where $N$ is the total number of n-grams.
We calculate the metrics for each Wikipedia article that has undergone primary filtering and normalise the scores across articles to be between 0 and 1. Following this, we sum the values for each metric within the respective class for each article to obtain the total score per class per article. This can be formalised as follows:
$S_{ij}=\sum_{k=1}^{K}\frac{M_{ijk}-M_{min}}{M_{max}-M_{min}}$ (1)
where: $S_{ij}$ is the total score for class $i$ in article $j$, $M_{ijk}$ is the value of metric $k$ in class $i$ for article $j$, $M_{min}$ is the minimum possible score for metric $k$, $M_{max}$ is the maximum possible score for metric $k$, and $K$ is the total number of metrics.
An important aspect of heuristic filters is that they rely on fixed thresholds in order to discriminate between high and low quality data. In much previous work, thresholds are typically user-defined and based on background assumptions of how ‘quality’ data should be distributed. However, doing so is not possible in our case, as such assumptions are often language-specific and cannot reliably generalise in a cross-lingual fashion. Thus, we implement an automatic thresholding method that is language-agnostic and distribution-specific. We determine thresholds for filtering Wikipedia data by comparing kernel density estimates of the lowest values and a random sample of values within each metric’s distribution. The procedure is detailed in Appendix C. Doing so allows us to dynamically set thresholds across metrics and Wikipedias. This, we find, filters fewer documents than the English-based Gopher thresholds across many languages.
4 Wikipedia Quality Tiers
Figure 2: Division of Wikipedia based on a clustering on fraction of documents removed and fraction of characters removed from global filtering. The blue cluster corresponds to Tier 1, orange to Tier 2, green to Tier 3 and red to Tier 4. Each point is weighed by the article count of the unfiltered Wikipedia.
Applying the primary filtering step to Wikipedia affects some languages more than others. Thus, we categorise non-English Wikipedias into 4 tiers: Tier 1, Tier 2, Tier 3 and Tier 4 based on how much of the Wikipedia is removed during primary filtering. We perform $k$-means clustering for all the languages where each language is represented by the percentage of documents remaining and the percentage of characters remaining in the Wikipedia after primary filtering. This is done to decorrelate quality and size. Based on several iterations of the clustering algorithm and manual inspections, we arrive at the categorisation illustrated in Figure 2. A list of all Wikipedias and their respective quality tiers is in Appendix D.
To put our categorisations into perspective, we compare them to the taxonomy of languages introduced by Joshi et al. (2020). They divide the world’s languages into 6 categories based on the amount of resources and labelled and unlabelled data (including Wikipedia) available: the Left Behinds, the Scraping Bys, the Hopefuls, the Rising Stars, the Underdogs and the Winners. A comparison of these categories and our quality tiers is in Figure 3.
Figure 3: A confusion matrix comparing the Joshi et al. (2020) and the Wikipedia quality tiers.
The Joshi et al. (2020) taxonomy has a clear correlation with the size of the Wikipedia and its categorisation, making this discussion a comparison of size and quality. We find that all Winners, unsurprisingly, end up in Tier 1 and Tier 2. However, most Scraping Bys and Left Behinds do as well, indicating that many languages, despite the shortage of NLP resources and smaller Wikipedias, have dedicated communities that are involved in the creation of good quality Wikipedias. There are also many Scraping Bys and Left Behinds that are in Tier 3 and Tier 4, such as Yorùbá and Twi, with serious quality issues that can be harmful to low-resource NLP. Lastly, we also observe some Underdogs and Rising Stars in Tier 3 (Swedish, Vietnamese, Serbian and Basque) and Tier 4 (Cebuano) categories, demonstrating that more availability of resources does not guarantee good quality. Thus, the NLP developer must be aware of such quality issues when using Wikipedia.
Heuristics as a Measure of Quality
The distribution of each Wikipedia for heuristic metrics defined in §3.2 and listed in Table 1 can be an indication of the general quality of the articles in that Wikipedia. The distributions of 8 Wikipedias, 2 from each tier on metrics from classes absolute, relative and entropy, defined in Table 1, are in Figure 4.
Figure 4: Distribution of select Wikipedias from Tier 1 (blue), Tier 2 (orange), Tier 3 (green) and Tier 4 (red) on each class of heuristic metrics.
The absolute distribution shows that there is a discrepancy in the distribution of characters to articles in Wikipedia, especially for the bottom three quality tiers. Most Wikipedias have a large head of very low scoring articles (between 0 and 0.1 out of 3) and a long tail of higher scoring articles. Yorùbá, for example, has more than 50% of its character count in just 10% of articles. The relative distribution also echoes this. Most Wikipedias have a spike at 2.0, the maximum possible score for that class, assigned to articles that do not have repeating words or trigrams. This typically correlates to very short, often one-line articles. Thus, the relative class favours shorter articles, with very short articles scoring the highest, and the rest of the Wikipedia following a more normal distribution. The absolute class, on the other hand, favours longer articles. entropy metrics favour a more normal distribution for higher quality Wikipedias, in Tier 1, and the bottom tiers with lower quality articles are more skewed.
5 Downstream Evaluation
In this section we evaluate the practical implications of applying the data filters from §3 for languages in the different tiers from §4. We evaluate the effect of data pruning in two scenarios: monolingual pretraining and language adaption of multilingual models.
Figure 5: The experimental pipeline
5.1 Monolingual Pretraining
To isolate the impact of data quality on downstream performance, we pretrain monolingual models, with a scaled down DeBERTa He et al. (2021) architecture on 51 Wikipedias from the 4 tiers. These languages are selected based on the availability of downstream evaluation tasks. Training small monolingual models minimises the risk of confounding factors affecting the results, while also being very efficient to train.
We study the effect of data pruning by pretraining models on three versions of Wikipedia: (1) the entire Wikipedia of the language without any filtering (raw_wiki), (2) filtered Wikipedia using the primary filters (+primary), and (3) further pruned using the heuristic filters (+heuristic), as described in §3. The full experimental pipeline is illustrated in Figure 5.
To understand the saliency of each class of heuristic filters detailed in §3.2, we sort the articles based on their total score and split them into high and low quality partitions such that both partitions have a roughly equal number of characters. We then pretrain monolingual models over these partitions for each class. This ensures that any differences in performance can be attributed to data quality and not size.
For every language, we pretrain 9 monolingual models. All models are subsequently fine-tuned on MasakhaNER 2.0 Adelani et al. (2022), MasakhaPOS Dione et al. (2023), MasakhaNEWS Adelani et al. (2023), AfriSenti-Twitter Muhammad et al. (2023) and SIB200 Adelani et al. (2024), based on availability for every language. Details about model and tokeniser configurations and the entire training regime are in Appendix A.
5.1.1 Effect of Data Quality Filtering
Comparing the effect of data quality filtering across various downstream tasks reveals that primary filtering generally is not detrimental to accuracy. Specific results per language and per task may vary. MasakhaNER (Table 2) on average has similar performance for no filtering and primary filtering, but has a slight performance drop with heuristic filtering. MasakhaNEWS (Table 3), on the other hand, has a stronger preference for primary filtering. AfriSenti-Twitter (Table 4) has the same average performance with or without any filtering.
We note that these averages are not entirely representative of trends with individual Wikipedias. For example with MasakhaNEWS, while there are minimal differences across the partitions for most languages, om and ti get a substantial performance boost from primary filtering. Conversely, rn performance after heuristic filtering is much worse than no filtering at all.
For AfriSenti-Twitter, although the average scores across partitions are the same, there is never a performance decline with +primary, except for ary. In fact, am has a 6% performance boost from no filtering to heuristic filtering. The ary performance drop can be explained by the fact that the ary subset in AfriSenti-Twitter contains script switching between Arabic and Latin. None of the models, except raw_wiki, is able to tokenise Latin characters due to the script filtering step which removes all Latin characters from the vocabulary. This connects to the discussion in §2.3 about what data quality actually means (more in §6).
MasakhaNER 2
Lang raw_wiki +primary +heuristic
pcm 0.77±0.00 0.76±0.00 0.76±0.00
ha 0.82±0.00 0.81±0.00 0.81±0.00
ig 0.85±0.00 0.85±0.00 0.84±0.00
sn 0.92±0.00 0.92±0.00 0.92±0.00
rw 0.76±0.05 0.79±0.01 0.77±0.05
tw 0.76±0.00 0.76±0.01 0.74±0.00
wo 0.75±0.01 0.73±0.01 0.72±0.01
ny 0.87±0.00 0.87±0.00 0.86±0.00
ee 0.81±0.00 0.81±0.01 0.79±0.00
bm 0.74±0.01 0.74±0.01 0.72±0.01
fon 0.74±0.01 0.74±0.00 0.76±0.01
sw 0.90±0.00 0.90±0.00 0.90±0.00
zu 0.83±0.01 0.82±0.01 0.81±0.01
lg 0.84±0.00 0.85±0.00 0.83±0.00
tn 0.80±0.01 0.80±0.00 0.79±0.01
yo 0.84±0.00 0.84±0.00 0.83±0.00
Avg 0.81 0.81 0.80
Table 2: Fine-tuning results on MasakhaNER 2. Results are averaged over 5 runs, and languages are grouped by tier.
MasakhaNEWS
Lang raw_wiki +primary +heuristic
pcm 0.90±0.04 0.91±0.00 0.87±0.04
ha 0.88±0.01 0.86±0.01 0.86±0.01
ig 0.85±0.01 0.85±0.01 0.84±0.01
sn 0.89±0.01 0.90±0.01 0.89±0.01
so 0.71±0.01 0.71±0.01 0.70±0.01
ln 0.78±0.01 0.76±0.01 0.76±0.01
om 0.67±0.30 0.82±0.01 0.81±0.30
ti 0.60±0.04 0.65±0.02 0.62±0.04
sw 0.81±0.01 0.82±0.00 0.82±0.01
am 0.84±0.04 0.87±0.01 0.85±0.04
lg 0.86±0.01 0.86±0.02 0.84±0.01
rn 0.75±0.01 0.76±0.02 0.71±0.01
yo 0.84±0.03 0.86±0.01 0.86±0.03
Avg 0.80 0.82 0.80
Table 3: Fine-tuning results on MasakhaNEWS. Results are averaged over 5 runs, and languages are grouped by tier.
AfriSenti-Twitter
Lang raw_wiki +primary +heuristic
pcm 0.61±0.01 0.61±0.01 0.61±0.01
ha 0.72±0.01 0.72±0.01 0.73±0.01
ig 0.74±0.00 0.75±0.00 0.75±0.00
rw 0.60±0.04 0.61±0.01 0.59±0.04
tw 0.62±0.01 0.64±0.01 0.64±0.01
ts 0.48±0.04 0.49±0.03 0.50±0.04
sw 0.55±0.03 0.59±0.02 0.56±0.03
am 0.40±0.07 0.41±0.04 0.46±0.07
ary 0.46±0.01 0.40±0.01 0.40±0.01
yo 0.68±0.01 0.68±0.01 0.68±0.01
Avg 0.59 0.59 0.59
Table 4: Fine-tuning results on AfriSenti-Twitter. Results are averaged over 5 runs, and languages are grouped by tier.
With SIB200 we compare the effectiveness of data quality filtering across the tiers defined in §4. Table 5 contains the results grouped by tier. On average, languages in Tier 1 have similar performance with raw_wiki and +primary, but experience a slight drop with +heuristic, a trend also seen in MasakhaNER and MasakhaNEWS. More languages in Tier 2 have the best performance with +heuristic, even though on average results are similar across partitions. On average, Tier 3 has better performance with raw_wiki. Notably, lij and ace show a 7% and 10% performance drop with heuristic filtering respectively, impacting the overall average. However, more Wikipedias achieve their highest scores with filtering than when trained on raw_wiki (5 with +primary, 5 with +heuristic and 4 with raw_wiki). For the Tier 4 Wikipedias, while the average does not demonstrate a strong preference for one partition over another, individual languages differ a lot. Cebuano, for instance, benefits substantially from primary filtering, with a 12% performance gain. It can be attributed to the deduplication step that removes most bot generated articles.
SIB200
Lang raw_wiki +primary +heuristic
Tier 1
ti 0.38±0.05 0.38±0.01 0.38±0.05
ha 0.77±0.02 0.78±0.02 0.75±0.02
ts 0.61±0.03 0.60±0.01 0.56±0.03
ig 0.76±0.01 0.77±0.02 0.74±0.01
sn 0.66±0.03 0.62±0.03 0.61±0.03
so 0.68±0.01 0.73±0.02 0.70±0.01
rw 0.77±0.01 0.77±0.01 0.73±0.01
tw 0.65±0.10 0.65±0.02 0.67±0.10
ln 0.59±0.02 0.57±0.02 0.56±0.02
wo 0.50±0.02 0.52±0.02 0.45±0.02
om 0.59±0.01 0.62±0.01 0.56±0.01
ny 0.64±0.02 0.60±0.01 0.64±0.02
ee 0.59±0.03 0.61±0.03 0.57±0.03
bm 0.45±0.02 0.48±0.02 0.47±0.02
fon 0.48±0.03 0.45±0.04 0.48±0.03
Avg 0.61 0.61 0.59
Tier 2
ro 0.82±0.03 0.80±0.02 0.81±0.03
id 0.83±0.01 0.85±0.01 0.85±0.01
hu 0.82±0.01 0.86±0.01 0.82±0.01
nl 0.82±0.03 0.80±0.03 0.82±0.03
uk 0.85±0.01 0.81±0.02 0.84±0.01
awa 0.54±0.03 0.52±0.03 0.54±0.03
ks 0.42±0.02 0.45±0.02 0.42±0.02
ckb 0.80±0.01 0.82±0.01 0.82±0.01
rn 0.51±0.03 0.50±0.02 0.51±0.03
sw 0.79±0.02 0.81±0.02 0.83±0.02
am 0.66±0.01 0.65±0.02 0.63±0.01
zu 0.73±0.01 0.70±0.02 0.69±0.01
tn 0.64±0.02 0.61±0.01 0.64±0.02
ary 0.59±0.00 0.60±0.01 0.61±0.00
lg 0.55±0.02 0.55±0.02 0.56±0.02
Avg 0.69 0.69 0.69
Tier 3
lij 0.61±0.01 0.59±0.02 0.54±0.01
ht 0.70±0.04 0.69±0.01 0.66±0.04
te 0.80±0.02 0.79±0.02 0.81±0.02
mr 0.74±0.04 0.78±0.02 0.76±0.04
kk 0.81±0.02 0.83±0.02 0.80±0.02
yo 0.70±0.02 0.69±0.02 0.70±0.02
ace 0.27±0.02 0.15±0.03 0.17±0.02
shn 0.74±0.02 0.72±0.02 0.75±0.02
mai 0.67±0.01 0.68±0.02 0.68±0.01
ban 0.65±0.01 0.66±0.02 0.68±0.01
vi 0.82±0.01 0.83±0.00 0.81±0.01
nso 0.50±0.02 0.53±0.01 0.46±0.02
Avg 0.67 0.66 0.65
Tier 4
tt 0.79±0.01 0.75±0.01 0.76±0.01
ceb 0.39±0.03 0.51±0.03 0.50±0.03
war 0.48±0.04 0.49±0.04 0.50±0.04
pag 0.45±0.03 0.48±0.01 0.46±0.03
mi 0.55±0.02 0.52±0.02 0.52±0.02
mni 0.27±0.04 0.21±0.01 0.21±0.04
azb 0.53±0.05 0.53±0.02 0.51±0.05
min 0.65±0.03 0.60±0.03 0.53±0.03
Avg 0.51 0.51 0.50
Table 5: Fine-tuning results on SIB200 for tiny DeBERTa models. Results are averaged over 5 runs.
5.1.2 Saliency of Heuristic Filters
Results for models trained on low and high quality partitions from each class of heuristic filters are illustrated in Figure 6. Scores for each model and task are in Appendix B. Notably, we find that models pretrained on essentially half the data of the +primary filtering and raw_wiki models perform just as well for most tasks. Recall that with models in Tables 2, 3, 4 and 5 performance with filtering is on par with performance without. This indicates that we do not need large amounts of data to achieve competitive performance on downstream tasks, with quality mattering more than quantity during pretraining.
Comparing the performance difference between each heuristic filter, there is a signal from the distribution of the pretraining data to downstream performance, but it is not consistent. Thus, we cannot draw strong conclusions about the efficacy of each individual heuristic filter. There is little variance across models, across tasks, which can also explain why the heuristic filtering step also has little effect on performance.
Figure 6: Fine-tuning results for models trained on the absolute_low (blue), absolute_high (orange), relative_low (green), relative_high (red), entropy_low (purple), and entropy_high (brown) partitions.
5.2 Language adaptation
While the tiny models allow us to fully control the experimental variables, we also test the effect of our filtering methods in a more practical scenario. Language adaption of existing multilingual models to an unseen language is a common practice for low-resource languages Pfeiffer et al. (2020, 2022). To test this scenario, we follow the experimental protocol from Alabi et al. (2022) in training AfroXLM-R and test the efficacy of filtering Wikipedia by performing language adaption on XLM-R base Conneau et al. (2020) on unseen languages from our experimental cohort. We use the original XLM-R tokeniser and further pretrain XLM-R on Wikipedia with raw_wiki, +primary, and +heuristic. We fine-tune each model on SIB200. Hyperparameter settings are listed in Appendix A.
XLMR Continued Pretraining
Lang baseline raw_wiki +primary +heuristic
ti 0.59±0.04 0.53±0.21 0.56±0.02 0.60±0.21
fon 0.43±0.20 0.60±0.04 0.56±0.02 0.57±0.04
ig 0.52±0.20 0.79±0.02 0.79±0.03 0.75±0.02
ee 0.64±0.04 0.66±0.03 0.64±0.03 0.64±0.03
ny 0.64±0.01 0.68±0.03 0.65±0.02 0.68±0.03
wo 0.60±0.03 0.65±0.01 0.64±0.01 0.65±0.01
ln 0.66±0.01 0.70±0.01 0.69±0.01 0.68±0.01
tw 0.66±0.02 0.72±0.02 0.70±0.03 0.70±0.02
rw 0.56±0.03 0.69±0.02 0.67±0.02 0.63±0.02
sn 0.54±0.03 0.59±0.02 0.60±0.03 0.60±0.02
bm 0.43±0.22 0.60±0.05 0.55±0.03 0.54±0.05
ts 0.58±0.05 0.63±0.04 0.60±0.05 0.59±0.04
Avg 0.57 0.65 0.64 0.64
tn 0.56±0.02 0.62±0.03 0.60±0.04 0.55±0.03
ks 0.62±0.22 0.68±0.02 0.68±0.02 0.68±0.02
rn 0.54±0.04 0.55±0.06 0.58±0.02 0.57±0.06
lg 0.40±0.17 0.58±0.03 0.58±0.03 0.55±0.03
zu 0.58±0.01 0.64±0.03 0.68±0.02 0.65±0.03
awa 0.80±0.04 0.81±0.02 0.83±0.03 0.80±0.02
Avg 0.59 0.65 0.66 0.63
ht 0.63±0.03 0.73±0.03 0.73±0.02 0.72±0.03
nso 0.55±0.02 0.59±0.04 0.63±0.02 0.57±0.04
lij 0.77±0.02 0.81±0.03 0.78±0.03 0.80±0.03
ace 0.41±0.09 0.45±0.13 0.53±0.06 0.51±0.13
shn 0.33±0.16 0.61±0.03 0.60±0.04 0.62±0.03
mai 0.82±0.02 0.83±0.02 0.82±0.02 0.85±0.02
ban 0.79±0.01 0.82±0.02 0.83±0.01 0.83±0.02
yo 0.56±0.02 0.63±0.04 0.66±0.02 0.62±0.04
Avg 0.61 0.68 0.70 0.69
mi 0.59±0.12 0.62±0.02 0.63±0.03 0.59±0.02
mni 0.54±0.05 0.38±0.11 0.45±0.03 0.45±0.11
pag 0.77±0.02 0.79±0.02 0.77±0.05 0.78±0.02
Avg 0.59 0.66 0.66 0.65
Table 6: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on SIB200. Results are averaged over 5 runs, and languages are grouped by tier.
Results are in Table 6, with the trends being slightly different from the trends seen with the monolingual models in Table 5. While all languages (except mni) benefit from language adaption, the effect of quality filtering is dependent upon the underlying data distribution. Tier 2 and Tier 3 benefit the most from primary filtering, with ace achieving an 8% increase. Tier 1 and Tier 4, on the other hand, do not see much performance improvement with filtering on average. Notably, mni suffers a staggering 16% performance drop after language adaption, raising many questions about the quality of that Wikipedia.
6 Discussion
6.1 Effect of Quality Filtering
We look for a correlation between data pruning and accuracy with a Spearman’s correlation test between the amount of data removed in the primary filtering step, and the accuracy difference between raw_wiki and +primary. There is no significant correlation, which means that data quality may not have a direct significant impact on performance. Findings from van Noord et al. (2024) and Artetxe et al. (2022) also echo this result. This means that there is no significant decrease in performance when removing much of the data. With results in Figure 6 we saw a reduction in data size by 50% with little impact on performance. Thus, we recommend data quality filtering not as a means for improving performance, but as a means for efficient training and achieving good performance with less compute.
Overall, the primary filtering step with script filtering and deduplication is beneficial for most Wikipedias. Further heuristic filtering, on the other hand, yields mixed results. This can be dependent upon the underlying distribution of the Wikipedia, since Wikipedias in Tier 2 and Tier 3 can benefit more from quality filtering than those in Tier 1, which already have good quality data distributions, and in Tier 4 with subpar data distributions (see Figure 4) that cannot be helped by filtering data any further.
6.2 Quality versus Size
Here we highlight the benefits of data quality over both data quantity and model size. While the discussions so far have mainly been centred around monolingual data quality, a comparison of results from Tables 5 and 6 highlight another salient aspect. When compared with the XLM-R model adapted to the same data (for e.g., monolingual raw_wiki versus XLM-R+raw_wiki), 23 monolingual models either perform on par or outperform XLM-R with a fraction of the data size and capacity. Our results for AfriSenti-Twitter are competitive with those reported by Muhammad et al. (2023) on larger multilingual models. This indicates that a smaller concentration of monolingual data trained with fewer parameters can yield better performance than a large multilingual corpus that includes the same data. Thus, low-resource languages benefit from smaller, more efficient models trained on a fraction of the compute. This is also seen with the models in Figure 6, which achieve competitive performance to models trained on the entire Wikipedia with essentially half the data.
6.3 Data Quality versus Domain
As discussed in §2.3, data quality is subjective and so domain may be a more useful consideration. We see a reflection of this when comparing the effect of data quality across tasks in Tables 2, 3 and 4. MasakhaNER favours the seemingly lower quality, unfiltered raw_wiki distribution. This may be due to the encyclopedic domain of Wikipedia lending itself very well to good performance on a task like NER. Quality filtering can remove the one-liner articles that describe named entities which can be beneficial for this task. Performance on classification tasks, on the other hand, benefits from the data quality filtering.
Thus, we push for a shift in perspective from seeking general definitions of data quality to domain and task specific investigations of what data quality means.
6.4 Data Quality versus Language
A push away from general definitions of data quality also leads us towards considering the effect of data quality on individual Wikipedias. The results discussed in this study on average remain consistent, but zooming in on individual languages reveals the large benefits of this process for some languages. We touched upon the results for specific languages in §5.1.1, showing that averages may not entirely reflect trends with individual Wikipedias. Therefore, we emphasize that there there is no one-size-fits-all data filtering approach. Data quality decisions for a corpus like Wikipedia containing multiple languages, even though they are all within the same domain, should be made on a per-language basis. This approach can be made more principled by leveraging typological features and databases to create more language-specific data filtering pipelines, which is a promising direction for future work.
7 Conclusion
With this study, we challenge the assumption of Wikipedia as a high quality resource in multilingual NLP, and provide a guide to using Wikipedia as a pretraining resource. We show that much of Wikipedia contains duplicate articles, foreign scripts, etc., which can be filtered out as a means to achieve good performance with less compute. For some languages, it can even lead to performance improvement. Overall, we advocate towards a shift in perspective from seeking general definitions of data quality to approaching quality issues from a domain or language specific lens.
8 Acknowledgements
For KT, AK, WP and MDL the computational resources and services used were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government - department EWI. AK was partially funded by a Google Research Award. WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. EP, HL and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF210454). We would also like to thank Isaac Caswell, Colin Cherry, and Markus Freitag at Google Translate Research for their insightful reviews on the initial draft.
Appendix A Pretraining Details
Hyperparameters for the pretraining and fine-tuning of monolingual models are in Table 7. We use the SentencePiece package Kudo and Richardson (2018) to train monolingual unigram tokenisers for each language. We train two tokenisers per language: one on the raw_wiki with no filtering, which is used to train the raw_wiki models; the second one on the data remaining after script filtering and deduplication, which we use to train all other models. We use an approximation with Zipf’s law Zipf (1949) on each Wikipedia to estimate the vocabulary size for each model, mentioned in Table 7. Hyperparameters for XLM-R language adaption are also in Table 7. For both cases, we checkpoint and use the best model on a left-out validation set to fine-tune after pretraining.
Tiny DeBERTa Hyperparameters Estimated Vocabulary Size
Pretraining Lang Vocab Lang Vocab
Model Type deberta ace 3505 mr 14708
Hidden Size 312 am 5234 nl 60011
Intermediate Size 1200 ary 4502 nso 2801
Attention Heads 12 awa 2190 ny 2472
Hidden Layers 4 azb 13996 om 3436
Hidden Act gelu ban 6624 pag 2170
Max Length 128 bm 1485 pcm 2595
Task MLM ceb 73124 rn 1446
Mask Probability 0.4 ckb 11451 ro 36329
Padding Strategy longest ee 1770 rw 5725
Grad. Accumulation Steps 4 fon 1426 shn 5784
Batch size 8 and 32 ha 13126 sn 5227
Learning Rate 1e-3 ht 10836 so 6117
Epochs 100 hu 46205 sw 12529
Finetuning id 41725 te 23175
Max Length 128 ig 11609 ti 1150
Epochs 20 kk 22496 tn 3783
Batch Size 8 ks 2248 ts 1817
Learning Rate 5e-5 lg 4528 tt 26047
Padding Strategy max_length lij 5462 tw 4745
XLMR Hyperparameters ln 2509 uk 62626
Language Adaption mai 4793 vi 43806
Model Type xlm-roberta mi 3451 war 25803
Task MLM min 14783 wo 3384
Max Length 512 mni 3351 yo 6128
Mask Probability 0.15 zu 4439
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding strategy Longest
Grad. Accumulation Steps 1
Finetuning
Max Length 512
Epochs 20
Batch Size 32
Learning Rate 5.00e-5
Padding Strategy max_length
Table 7: (1) Hyperparameters for monolingual pretraining and fine-tuning. While we used 100 epochs as a default, for Wikipedias that are too large (nl, ro, id, hu, uk, vi, ceb, war, tt, azb, min) we pretrain only for 2 epochs. Similarly, while we used a batch size of 8 for all models, for the larger Wikipedias we use a batch size of 32. (2) Hyperparameters for XLM-R language adaptation (based on Alabi et al. (2022)) and fine-tuning. (3) Estimated vocabulary size used to train monolingual tokenisers for each Wikipedia in our experimental pipeline.
Appendix B Results from partitioned models
Results for the partitioned models fine-tuned on SIB200 are in Table 8, MasakhaNER, MasakhaNEWS and AfriSenti-Twitter are in Table 9. Scores from this table were used to create the boxplot in Figure 6.
SIB200
Lang absolute relative entropy
low high low high low high
ha 0.77 0.77 0.75 0.71 0.72 0.77
ig 0.76 0.76 0.74 0.75 0.73 0.77
sn 0.59 0.64 0.62 0.62 0.54 0.66
so 0.71 0.65 0.64 0.69 0.72 0.63
rw 0.68 0.72 0.72 0.72 0.71 0.74
tw 0.67 0.64 0.66 0.65 0.64 0.67
ln 0.54 0.56 0.52 0.54 0.53 0.56
wo 0.50 0.41 0.42 0.51 0.45 0.41
om 0.54 0.50 0.54 0.58 0.57 0.56
ny 0.63 0.59 0.63 0.55 0.56 0.61
ee 0.57 0.57 0.62 0.59 0.58 0.59
bm 0.38 0.42 0.49 0.39 0.46 0.39
ts 0.49 0.54 0.46 0.52 0.52 0.50
fon 0.46 0.45 0.43 0.44 0.45 0.46
ti 0.38 0.28 0.34 0.36 0.37 0.34
sw 0.79 0.78 0.81 0.79 0.73 0.82
am 0.64 0.57 0.61 0.56 0.61 0.62
zu 0.63 0.66 0.68 0.58 0.64 0.68
ary 0.63 0.57 0.57 0.56 0.57 0.60
lg 0.47 0.56 0.52 0.52 0.51 0.51
tn 0.60 0.61 0.58 0.59 0.58 0.60
rn 0.45 0.48 0.46 0.52 0.48 0.51
ckb 0.80 0.82 0.81 0.81 0.81 0.81
ks 0.41 0.43 0.46 0.38 0.43 0.40
awa 0.53 0.46 0.53 0.52 0.51 0.52
ro 0.79 0.80 0.79 0.77 0.83 0.79
id 0.85 0.79 0.83 0.85 0.80 0.85
hu 0.81 0.83 0.82 0.85 0.81 0.85
nl 0.80 0.82 0.86 0.81 0.82 0.83
uk 0.82 0.83 0.81 0.83 0.84 0.83
yo 0.66 0.64 0.67 0.70 0.63 0.56
kk 0.81 0.82 0.82 0.82 0.83 0.80
mr 0.78 0.77 0.77 0.75 0.73 0.78
te 0.80 0.81 0.82 0.80 0.83 0.86
ht 0.68 0.66 0.71 0.69 0.65 0.70
nso 0.41 0.54 0.57 0.48 0.41 0.48
lij 0.55 0.55 0.46 0.57 0.54 0.52
ace 0.18 0.17 0.16 0.19 0.22 0.23
shn 0.71 0.74 0.74 0.69 0.73 0.74
mai 0.64 0.65 0.61 0.61 0.65 0.66
ban 0.59 0.66 0.64 0.68 0.64 0.66
mni 0.24 0.16 0.21 0.22 0.18 0.24
mi 0.56 0.54 0.55 0.51 0.53 0.52
pag 0.44 0.39 0.39 0.40 0.44 0.47
vi 0.82 0.84 0.82 0.83 0.83 0.84
war 0.46 0.54 0.56 0.47 0.47 0.59
tt 0.74 0.72 0.77 0.69 0.73 0.69
azb 0.49 0.53 0.48 0.42 0.50 0.47
ceb 0.46 0.50 0.49 0.47 0.52 0.53
min 0.54 0.51 0.47 0.48 0.49 0.49
Table 8: Fine-tuning results for monolingual models on SIB200. Languages are segmented by tier.
Masakhane
Lang absolute relative entropy
low high low high low high
MasakhaNER 2
pcm 0.75 0.74 0.74 0.75 0.74 0.74
ha 0.80 0.79 0.81 0.81 0.82 0.82
ig 0.85 0.84 0.84 0.85 0.84 0.84
sn 0.90 0.91 0.90 0.91 0.91 0.92
rw 0.77 0.76 0.78 0.78 0.76 0.76
tw 0.74 0.74 0.74 0.74 0.74 0.74
wo 0.74 0.69 0.72 0.74 0.73 0.71
ny 0.86 0.86 0.87 0.86 0.86 0.86
ee 0.79 0.78 0.79 0.79 0.79 0.79
bm 0.72 0.68 0.72 0.72 0.71 0.72
fon 0.73 0.71 0.71 0.71 0.74 0.74
sw 0.90 0.90 0.90 0.90 0.90 0.90
zu 0.82 0.80 0.82 0.80 0.80 0.82
lg 0.82 0.83 0.82 0.84 0.83 0.83
tn 0.77 0.78 0.77 0.78 0.76 0.78
yo 0.83 0.84 0.83 0.84 0.83 0.84
MasakhaNEWS
pcm 0.88 0.90 0.90 0.88 0.89 0.89
ha 0.86 0.88 0.87 0.85 0.86 0.87
ig 0.86 0.84 0.86 0.83 0.86 0.83
sn 0.90 0.89 0.88 0.89 0.87 0.90
so 0.69 0.68 0.70 0.69 0.68 0.69
ln 0.77 0.79 0.81 0.81 0.77 0.79
om 0.79 0.79 0.80 0.80 0.78 0.80
ti 0.63 0.59 0.61 0.63 0.62 0.60
sw 0.80 0.81 0.80 0.82 0.83 0.80
am 0.86 0.86 0.86 0.86 0.85 0.86
lg 0.86 0.86 0.88 0.85 0.88 0.86
rn 0.71 0.72 0.69 0.72 0.70 0.69
yo 0.84 0.85 0.85 0.87 0.86 0.86
AfriSenti-Twitter
pcm 0.60 0.60 0.59 0.60 0.60 0.60
ha 0.74 0.73 0.73 0.72 0.70 0.73
ig 0.74 0.75 0.74 0.75 0.74 0.73
rw 0.59 0.59 0.59 0.59 0.59 0.60
tw 0.64 0.62 0.63 0.65 0.63 0.63
ts 0.48 0.50 0.54 0.50 0.53 0.50
sw 0.54 0.55 0.58 0.56 0.59 0.57
am 0.40 0.43 0.48 0.32 0.28 0.14
ary 0.40 0.40 0.39 0.42 0.40 0.41
yo 0.67 0.68 0.67 0.67 0.67 0.69
Table 9: Fine-tuning results for monolingual models on MasakhaNER, MasakhaNEWS and AfriSenti-Twitter. Languages are segmented by tier.
Appendix C Thresholding Details
We detail the procedure for creating thresholds for heuristic metrics as follows: for a given Wikipedia metric distribution (e.g., rps_doc_word_count), we determine a representative sample size based on the full dataset. Testing across various metrics and Wikipedias, we select a sample size of $n_{\mathrm{sample}}=0.05*n_{\mathrm{docs}}$, where $n_{\mathrm{docs}}$ is the total number of documents in the dataset. We then sort the full distribution in ascending order and take the first $n_{\mathrm{sample}}$ values as the value distribution, $D_{\mathrm{value}}$. This distribution reflects the lowest values in the distribution —essentially, the values we want to assess as potential outliers or benchmarks. Separately, we randomly sample $n_{\mathrm{sample}}$ values from the full distribution to create the sampled distribution, $D_{\mathrm{sample}}$. This random sample provides a baseline density that represents the overall data distribution rather than just extreme values. We estimate the kernel density (KDE) for both $D_{\mathrm{value}}$ and $D_{\mathrm{sample}}$, applying these KDEs over $n_{\mathrm{sample}}$ evenly spaced values in the range:
$\left[\min(D_{\mathrm{value}}),\,\max(D_{\mathrm{sample}})\right]$
The threshold is then identified as the point within this range where the difference between the two density estimates is maximized.
Appendix D Quality Tiers
Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier
ab Tier 3 crh Tier 3 ha Tier 1 lez Tier 1 pa Tier 1 su Tier 3
ace Tier 3 cs Tier 1 hak Tier 3 lfn Tier 1 pag Tier 4 sv Tier 3
ady Tier 1 csb Tier 1 haw Tier 3 lg Tier 2 pam Tier 2 sw Tier 2
af Tier 1 cu Tier 1 he Tier 1 li Tier 1 pap Tier 1 szl Tier 3
als Tier 1 cv Tier 4 hi Tier 2 lij Tier 3 pcd Tier 1 ta Tier 1
alt Tier 2 cy Tier 2 hif Tier 2 lld Tier 4 pcm Tier 1 tay Tier 2
am Tier 2 da Tier 1 hr Tier 1 lmo Tier 3 pdc Tier 1 tcy Tier 1
ami Tier 2 dag Tier 3 hsb Tier 2 ln Tier 1 pfl Tier 3 te Tier 3
an Tier 1 de Tier 1 ht Tier 3 lo Tier 2 pi Tier 4 tet Tier 1
ang Tier 1 din Tier 1 hu Tier 2 lt Tier 1 pih Tier 1 tg Tier 3
ar Tier 2 diq Tier 3 hy Tier 2 ltg Tier 1 pl Tier 1 th Tier 1
arc Tier 1 dsb Tier 1 hyw Tier 1 lv Tier 1 pms Tier 3 ti Tier 1
arz Tier 3 dty Tier 2 ia Tier 3 mad Tier 1 pnb Tier 2 tk Tier 2
as Tier 1 dv Tier 1 id Tier 2 mai Tier 3 pnt Tier 2 tl Tier 1
ast Tier 2 dz Tier 1 ie Tier 3 map-bms Tier 3 ps Tier 1 tly Tier 3
atj Tier 2 ee Tier 1 ig Tier 1 mdf Tier 3 pt Tier 1 tn Tier 2
av Tier 2 el Tier 1 ik Tier 1 mg Tier 3 pwn Tier 1 to Tier 1
awa Tier 2 eml Tier 3 ilo Tier 2 mhr Tier 2 qu Tier 2 tpi Tier 2
ay Tier 1 eo Tier 1 inh Tier 1 mi Tier 4 rm Tier 2 tr Tier 1
az Tier 2 es Tier 1 io Tier 2 min Tier 4 rmy Tier 1 trv Tier 1
azb Tier 4 et Tier 1 is Tier 1 mk Tier 2 rn Tier 2 ts Tier 1
ba Tier 2 eu Tier 3 it Tier 1 ml Tier 1 ro Tier 2 tt Tier 4
ban Tier 3 ext Tier 1 iu Tier 2 mn Tier 1 roa-rup Tier 1 tum Tier 3
bar Tier 1 fa Tier 2 ja Tier 1 mni Tier 4 roa-tara Tier 2 tw Tier 1
bat-smg Tier 3 fat Tier 1 jam Tier 1 mnw Tier 1 ru Tier 1 ty Tier 3
bcl Tier 2 ff Tier 1 jbo Tier 1 mr Tier 3 rue Tier 2 tyv Tier 1
be Tier 2 fi Tier 1 jv Tier 2 mrj Tier 2 rw Tier 1 udm Tier 2
be-x-old Tier 1 fiu-vro Tier 1 ka Tier 2 ms Tier 3 sa Tier 2 ug Tier 2
bg Tier 1 fj Tier 1 kaa Tier 1 mt Tier 1 sah Tier 1 uk Tier 2
bh Tier 2 fo Tier 1 kab Tier 1 mwl Tier 1 sat Tier 1 ur Tier 2
bi Tier 1 fon Tier 1 kbd Tier 1 my Tier 2 sc Tier 1 uz Tier 2
bjn Tier 3 fr Tier 1 kbp Tier 1 myv Tier 2 scn Tier 2 ve Tier 3
blk Tier 1 frp Tier 1 kcg Tier 1 mzn Tier 3 sco Tier 1 vec Tier 3
bm Tier 1 frr Tier 1 kg Tier 2 nah Tier 2 sd Tier 1 vep Tier 1
bn Tier 1 fur Tier 1 ki Tier 2 nap Tier 3 se Tier 2 vi Tier 3
bo Tier 2 fy Tier 1 kk Tier 3 nds Tier 1 sg Tier 1 vls Tier 1
bpy Tier 3 ga Tier 1 kl Tier 1 ne Tier 1 sh Tier 3 vo Tier 2
br Tier 1 gag Tier 2 km Tier 1 new Tier 4 shi Tier 2 wa Tier 1
bs Tier 2 gan Tier 3 kn Tier 1 nia Tier 2 shn Tier 3 war Tier 4
bug Tier 4 gcr Tier 1 ko Tier 1 nl Tier 2 si Tier 1 wo Tier 1
bxr Tier 1 gd Tier 1 koi Tier 3 nn Tier 1 simple Tier 1 wuu Tier 1
ca Tier 1 gl Tier 1 krc Tier 2 no Tier 1 sk Tier 2 xal Tier 3
cbk-zam Tier 3 glk Tier 3 ks Tier 2 nov Tier 2 sl Tier 1 xh Tier 2
cdo Tier 3 gn Tier 1 ksh Tier 1 nrm Tier 2 sm Tier 1 xmf Tier 2
ce Tier 4 gom Tier 1 ku Tier 3 nso Tier 3 sn Tier 1 yi Tier 1
ceb Tier 4 gor Tier 3 kv Tier 2 nv Tier 3 so Tier 1 yo Tier 3
ch Tier 2 got Tier 2 kw Tier 1 ny Tier 1 sq Tier 2 za Tier 2
chr Tier 3 gu Tier 3 ky Tier 2 oc Tier 2 sr Tier 3 zea Tier 2
chy Tier 1 guc Tier 1 la Tier 2 olo Tier 1 srn Tier 3 zh Tier 1
ckb Tier 2 gur Tier 1 lad Tier 1 om Tier 1 ss Tier 1 zh-classical Tier 1
co Tier 1 guw Tier 1 lb Tier 1 or Tier 2 st Tier 2 zh-min-nan Tier 4
cr Tier 3 gv Tier 1 lbe Tier 2 os Tier 2 stq Tier 1 zh-yue Tier 2
zu Tier 2
Table 10: Quality categorisations for all non-English Wikipedias.