### v2

# How Good is Your Wikipedia? Auditing Data Quality for Low-resource and Multilingual NLP

Abstract
Wikipedia’s perceived high quality and broad language coverage have established it as a fundamental resource in multilingual NLP. In the context of low-resource languages, however, these quality assumptions are increasingly being scrutinised. This paper critically examines the data quality of Wikipedia in a non-English setting by subjecting it to various quality filtering techniques, revealing widespread issues such as a high percentage of one-line articles and duplicate articles. We evaluate the downstream impact of quality filtering on Wikipedia and find that data quality pruning is an effective means for resource-efficient training without hurting performance, especially for low-resource languages. Moreover, we advocate for a shift in perspective from seeking a general definition of data quality towards a more language- and task-specific one. Ultimately, we aim for this study to serve as a guide to using Wikipedia for pretraining in a multilingual setting. We also introduce a package that implements our data cleaning, tokenization, and model training pipeline.
1 Introduction
**footnotetext: Equal contribution
Wikipedia is an invaluable resource in NLP, with much of its usefulness steming from its open-source nature through which a community of authors, editors, and administrators can collaborate on the documentation of an ever-changing world. Such a framework ensures that information contained therein is not only current, but also well-written, accurate, and relevant to any given subject. For these reasons, Wikipedia has been employed for a myriad of NLP applications, such as language model (LM) training and benchmarking (Merity et al., 2016), question-answering (Hewlett et al., 2016; Rajpurkar et al., 2016; Joshi et al., 2017) and knowledge-base creation (Vrandečić and Krötzsch, 2014; Lehmann et al., 2024), among many others. Wikipedia is also a staple resource in multilingual NLP, where it is commonly employed for pretraining multilingual LMs (Devlin et al., 2019; Conneau and Lample, 2019) and creating benchmarks for downstream tasks, such as named-entity recognition Rahimi et al. (2019) and machine translation Schwenk et al. (2021).
Recently, increasing attention to data quality has led to the adoption of various commonplace data filtering practices within NLP, whereby noisy web-scale data is distilled into a more salient signal for LM pretraining. Wikipedia has often been employed as a ‘high quality’ reference in this process (e.g., Wenzek et al., 2020), largely due to its encyclopaedic domain and culture of rigorous community moderation. However, the assumption of Wikipedia’s inherent ‘high quality’ — though generally reliable for English and select high-resource languages — is now increasingly being challenged in the multilingual setting Kreutzer et al. (2022a). Indeed, it has been found that Wikipedias for lower-resourced languages can contain unnatural, inorganic text written, often machine generated, and other noise, which can be incomprehensible to native speakers Alabi et al. (2020); Lent et al. (2024). As a result, poor data quality for such languages leads to degrading performance and untrustworthy models, ultimately providing unusable technologies for speakers Held et al. (2023); Nicholas and Bhatia (2023); Durmus et al. (2024).
In this study, we critically examine the extent of Wikipedia’s ‘high quality’ in a non-English setting. Specifically, we compare common techniques for data filtering and assess how they perform on Wikipedia, with an emphasis on low-resource settings. We empirically demonstrate a lack of apparent data quality, showing that many non-English Wikipedias contain foreign characters and duplicate articles, which we remove in a primary filtering step, followed by the application of corpus-driven heuristic quality filters. Downstream evaluation reveals that our quality filters preserve, sometimes improve, downstream model performance, and thus can be used as a means for more resource-efficient training. Ultimately, we argue for a shift towards task and language-specific approaches to data quality, as opposed to one-size-fits-all approaches thereof — especially when working with lower-resource languages. Overall, our contributions are as follows:
  1. 1.
    We present empirical evidence of quality issues across Wikipedia and show that standard data cleaning pipelines can be used to mitigate them.
  2. 2.
    We evaluate the effect of data quality filtering on downstream performance, particularly in a lower-resource setting.
  3. 3.
    We make recommendations for using Wikipedia as a resource for dataset creation and pretraining models.
  4. 4.
    We compile a Wikipedia-specific data filtering pipeline and release code that can be used to replicate all our experiments. [^1] https://github.com/akulmizev/texieve.git
With this, we hope to provide a meaningful contribution to the field’s larger ongoing debate about data quantity versus quality. As such, we aim for this study to serve as a starting guide for using Wikipedia as a data resource in language modelling and beyond.
2 Related Work
2.1 Data Filtering
Recent studies in large language modelling have demonstrated the benefits of filtering web-scale data Wenzek et al. (2020); Raffel et al. (2023); Penedo et al. (2024); 01.AI et al. (2024). In this context, data filtering typically involves a combination of techniques, ranging from basic document-level heuristics to corpus-level statistics that serve as quality proxies (Albalak et al., 2024). For example, the popular C4 dataset includes heuristics that remove documents containing code (e.g., ‘Javascript’), boiler plate legal policies (e.g., ‘Terms and Conditions’), or inappropriate terms, among other hand-crafted rules Raffel et al. (2023). Alongside heuristics, deduplication is another common filtering technique designed to remove duplicate documents from a given dataset, only keeping one copy thereof. A particularly popular algorithm is MinHash, due to its effectiveness in removing semantically identical documents, as well as its scalability across CPU nodes Penedo et al. (2024). MinHash has likewise been combined with downstream clustering-based deduplication to filter documents in a more aggressive manner Tirumala et al. (2023). It was also part of the data cleaning pipeline of the multilingual CulturaX 2024 dataset, in combination with language identification, URL-based filtering, and metric-based filtering with data-driven thresholds automatically identified. More sophisticated techniques have recently also been employed: for example, 01.AI et al. (2024) leverage learned filters (i.e., quality classifiers and cluster-based filters), which are designed to preserve documents resembling a ‘high quality’ reference corpus – typically Wikipedia Wenzek et al. (2020); Together Computer (2023).
2.2 Wikipedia in Multilingual NLP
Wikipedia’s widespread language coverage and reputation as a high-quality data trove has made it a vital resource in multilingual NLP applications. It has served as the backbone of many multilingual benchmark datasets, such as WikiAnn Pan et al. (2017), XQuAD Artetxe et al. (2020), MLQA Lewis et al. (2020), TyDiQA-GoldP Clark et al. (2020), and BUCC Zweigenbaum et al. (2017), as well as a common source of pretraining data for popular multilingual LMs, such as mBERT Devlin et al. (2019) and XLM Lample and Conneau (2019). Beyond this, it was notably employed as the training data for CCNet’s language identification classifier, which is primarily used for partitioning the popular CommonCrawl dataset along language boundaries. CommonCrawl serves as the basis for many contemporary open-source pretraining datasets, therefore the quality and performance of such a classifier is paramount in retaining relevant data.
As a consequence of its ubiquity in NLP, Wikipedia has recently attracted increased scrutiny for its data quality. For example, Kreutzer et al. (2022a) find that the Wikipedia-oriented WikiMatrix Schwenk et al. (2021) contains a number of deprecated, incorrect, or otherwise mislabelled language codes, among the 86 languages included Kreutzer et al. (2022a). Lignos et al. (2022) identify similar shortcomings for lower-resourced languages in WikiAnn and even Wikidata. [^2] https://wikidata.org Studies on selected low-resource languages likewise confirm such findings. For example, Alabi et al. (2020) observe that the Yorùbá and Twi Wikipedias are ostensibly written by non-native speakers and lack the necessary diacritics and variation to be relevant to speakers. Meanwhile, Lent et al. (2024) find that sentences from Wikipedias for Creole languages are often incomprehensible by speakers and require manual correction before being used to create benchmarks.
2.3 Data Quality versus Data Quantity
The notion of data quality is inherently subjective. Quantitatively, it is commonly measured by performance gains on a downstream NLP task. Qualitatively, it can be understood to signify utterances that are accepted as valid by proficient speakers of a language, or, alternatively, text samples that do not promote toxic or offensive sentiments Kreutzer et al. (2022b); van Noord et al. (2024). Such definitions can sometimes be at odds with one another, however. For example, Longpre et al. (2024) show that a smaller, carefully filtered dataset can improve performance on downstream tasks in contrast to a larger, unfiltered dataset, but that the resulting model becomes more likely to generate toxic content. Likewise, diminishing gains have been demonstrated from training on ever-increasing amounts of data Warstadt et al. (2023). For example, language models pretrained on C4 (745GB) marginally outperform those trained on Wikipedia (16GB) over benchmarks like GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and SQuAD (Rajpurkar et al., 2016). However, many of these benchmarks were sourced from Wikipedia itself, suggesting that quantity is less important than domain Raffel et al. (2023).
3 Is Wikipedia ‘High Quality’?
Albalak et al. (2024) outline a set of best practices for quality data selection and filtering. In this work, we examine the effects of their proposed model-agnostic filtering strategies, [^3] We exclude classifier-based and perplexity-based filtering strategies measuring quality relative to a selection of data. as applied to Wikipedia. Conceptually, we can recognize such filters as belonging to two distinct categories based on their underlying objective: 1) procedures for removing undesirable material from Wikipedia, such as foreign scripts and duplicate articles; and 2) heuristics that aim to enhance the existing data distribution by removing otherwise lower quality text according to a given metric (e.g. article length in characters). For the first category, we posit that if this step removes a substantial amount of text from a given Wikipedia dump, the removed content was not of high quality to begin with. We refer to this step as primary filtering. For the second, we aim to further refine Wikipedia articles via heuristic filters for even higher quality. Such a categorisation is also employed in the creation of MADLAD-400 Kudugunta et al. (2023), where text is deduplicated in a preliminary step, followed by language identification and then heuristic-based filters are applied.
3.1 Primary Filtering
Language Filtering
Applying an existing language identification model to Wikipedia is a challenge due to data contamination – most models supporting all of Wikipedia’s languages have generally also been trained on Wikipedia (see e.g. Kargaran et al., 2023). Thus, we employ script-based filtering as a proxy for language identification by filtering characters in the Wikipedia that are not in the officially documented script(s) [^4] For example, the Konkani Wikipedia features 3 scripts – Devanagari, Latin and Kannada. for the concerned language. [^5] https://en.wikipedia.org/wiki/List_of_Wikipedias Figure 1 demonstrates that 3.37% of characters in non-English Wikipedias are affected by this. Additionally, various Wikipedias contain full articles written in a different language, generally English. [^6] We acknowledge that this methodology would not be effective for such English articles in non-English Wikipedias written in the Latin script. However, this would require a language ID model that has not been trained on Wikipedia For example, 6% of Assamese Wikipedia consists of foreign characters, where numerous articles are written almost entirely in English Wilson (2019). The implications of this are inevitably made clear when training a tokeniser on this Wikipedia, where much of the vocabulary budget can be used up by foreign characters.
Figure 1: Proportion of articles and characters filtered from non-English Wikipedias with primary filtering.
Exact-match Deduplication
We employ a hashing function to identify and remove articles which have the exact same text, but a different unique ID or title. We find that some Wikipedias are strongly affected by this process — for example, the Yorùbá Wikipedia loses almost 60% of its articles — many of which contain a single token Ìtọḱasí (Reference). Typically, articles removed by exact-match deduplication are placeholders containing filler or templatic text and generally contain little usable information. Figure 1 shows that 1.01% of non-English Wikipedias articles are exact duplicates.
MinHash Deduplication
While exact-match is able to filter identical documents, it cannot account for partial duplicates — i.e., sets of documents containing mostly identical text, but varying in terms of single sentences, paragraphs, or even hyperlinks. For example, the Yorùbá articles for the Baluch people and the Baluchi language are largely identical, with the latter containing an additional language data panel. To capture such cases, we perform MinHash deduplication with locality sensitive hashing (LSH) and discount any article that returns a Jaccard similarity of 0.85. We find that a staggering 28.33% of all non-English Wikipedia articles (and 8.18% characters) are affected by this procedure, largely consisting of Wikipedias known to be majority bot generated (more in §4.3).
Heuristic-based Quality Filters
Class Name Calculation
Absolute length length in characters.
unique_trigrams total no. of unique trigrams.
unique_words total no. of unique words.
Relative frac_unique_trigrams unique trigrams / total trigrams
frac_unique_words unique words / total words
Entropy trigram_entropy $-\sum_{t\in T}\left(\frac{\text{count}(t)}{N}\right)\cdot\log_{2}\left(\frac{\text{count}(t)}{N}\right)$
unigram_entropy
Table 1: The quality heuristics applied to Wikipedia, calculated for each article. For entropy, $t$ is a unique n-gram in a set of $T$ unique document n-grams, where $N$ is the total number of n-grams.
3.2 Heuristic Filtering
Albalak et al. (2024) categorise a process in which document-level statistics are calculated in order to remove low quality data beyond a given threshold as heuristic filtering. Heuristics, in this sense, refer to simple metrics that aim to numerically characterise document content — for example, the number of characters contained in a given document. The red-pajama-data-v2 (Together Computer, 2023) dataset cleaning pipeline offers a collection of many such heuristics, culled from previous data curation efforts (Rae et al., 2021; Raffel et al., 2020; Longpre et al., 2024; Penedo et al., 2024). In theory, if Wikipedia is assumed to be a ‘high quality’ data source, it should be robust to these heuristics and not suffer from large-scale document removal.
To determine if this is the case, we explore natural language heuristics that are relevant to the encyclopedic nature of Wikipedia and easily applicable in a multilingual setting. [^7] For example, rps_doc_stop_word_fraction is not employed because it is difficult to principally obtain stopword lists for all the languages in Wikipedia. These are rps_doc_frac_unique_words (type-token ratio), rps_doc_unigram_entropy, and rps_doc_word_count – as collected by RedPajama. We also experiment with several adaptations to these metrics, such as utilising trigrams or characters as the basic unit of measurement (instead of whitespace-separated words). Naturally, some of these metrics demonstrate a very high correlation with each other; for example, absolute counts of characters, words, and trigrams typically exhibit a very high positive pairwise correlation across documents ($\rho>0.95$). As such, we decide to group all metrics into three classes: absolute, relative, and entropy. We list all metrics and their associated classes in Table 1.
We calculate the metrics for each Wikipedia article that has undergone primary filtering and normalise the scores across articles to be between 0 and 1. Following this, we sum the values for each metric within the respective class for each article to obtain the total score per class per article. This can be formalised as follows:
$S_{ij}=\sum_{k=1}^{K}\frac{M_{ijk}-M_{min}}{M_{max}-M_{min}}$ (1)
where: $S_{ij}$ is the total score for class $i$ in article $j$, $M_{ijk}$ is the value of metric $k$ in class $i$ for article $j$, $K$ is the total number of metrics, and $M_{min}$ and $M_{max}$ are the minimum and maximum possible score for metric $k$, respectively.
An important aspect of heuristic filters is that they rely on fixed thresholds in order to discriminate between high and low quality data. In much previous work, thresholds are typically user-defined and based on background assumptions of how ‘quality’ data should be distributed (Rae et al., 2021; Marion et al., 2023; Raffel et al., 2023; Penedo et al., 2024). However, doing so is not possible in our case, as such assumptions are often language-specific and cannot reliably generalise in a cross-lingual fashion. Thus, we implement an automatic thresholding method that is language-agnostic and distribution-specific. This procedure is inspired by the heuristic filter tuning process of (Penedo et al., 2024), who manually inspect metric histograms for high and low quality proxy datasets (locally and globally deduplicated common crawl snapshots), and select a threshold according to where values in the low quality histogram exceed those in its high quality counterpart. Though effective, this approach does not apply to our use case in that 1) we do not have access to low or high quality proxy datasets, and 2) we operate in a multilingual space, where separate thresholds must be selected per language. Instead, we work with the heuristic metric distribution directly, taking $n$ documents from the tail (or head) to serve as a sample of potential outliers. We compare this with a random sample of $n$ documents from the overall distribution by estimating the kernel density (KDE) of both samples, which is ultimately used for determining the threshold. More details on this procedure are detailed in Appendix B. The outlined method allows us to dynamically set thresholds across metrics and Wikipedias, which, we find, filters fewer documents than the English-based Gopher thresholds across many languages.
4 Characterizing Quality
In this section, we attempt to shed light on how our chosen filtering strategies relate to a generalized notion of dataset quality, as it pertains to Wikipedia.
4.1 Establishing Quality Tiers
In applying the primary filtering procedure to Wikipedia, we find that it affects certain languages much more than others. As such, we attempt to categorise all non-English Wikipedias into 4 tiers: Tier 1, Tier 2, Tier 3 and Tier 4 based on how much of the Wikipedia is removed during primary filtering. To do so, we perform $k$-means clustering, where each Wikipedia is represented by the percentage of documents and characters remaining after primary filtering. Based on several iterations of the clustering algorithm and manual inspections, we arrive at the taxonomy illustrated in Figure 2. A list of all Wikipedias and their respective quality tiers can be found in Table 16.
Figure 2: Division of Wikipedia based on a clustering on fraction of documents removed and fraction of characters removed from primary filtering. The blue cluster corresponds to Tier 1, orange to Tier 2, green to Tier 3 and red to Tier 4. Each point is weighed by the article count of the unfiltered Wikipedia.
Figure 3: A confusion matrix comparing the Joshi et al. (2020) and the Wikipedia quality tiers.
In order to put our quality taxonomy into perspective, we compare it to the taxonomy of languages introduced by Joshi et al. (2020), who divide the world’s languages into 6 categories based on the amount of labelled and unlabelled data (including Wikipedia) available for them. The categories they employ are as follows: the Left Behinds, the Scraping Bys, the Hopefuls, the Rising Stars, the Underdogs and the Winners. Figure 3 displays a comparison of these categories and our quality tiers.
Joshi et al. (2020)’s taxonomy correlates strongly with Wikipedia size, yet it yields limited alignment with our quality taxonomy. We interpret this as indicating that quality does not follow directly from resources. Indeed, while we observe that most Winners, unsurprisingly, end up in Tier 1, Arabic lands in Tier 2. Upon a cursory inspection, we find that 15% of the original Arabic Wikipedia is removed via MinHash deduplication, indicating that it contains a sizable volume of potentially templatic or bot-generated text. Going further, we find that 113 of the lowest resourced languages classified as Scraping Bys and Left Behinds (35% of Wikipedia in total) are likewise categorized as having Tier 1 Wikipedias, implying the involvement of dedicated moderation communities. Conversely, we also observe many Scraping Bys and Left Behinds languages falling into the Tier 3 and Tier 4 categories, such as Yorùbá and Twi, indicating that they contain serious quality issues that can be harmful to low-resource NLP. Lastly, we note that various higher-resourced Underdogs and Rising Stars languages are placed in Tier 3 (Swedish, Vietnamese, Serbian and Basque) and Tier 4 (Cebuano). Such Wikipedias tend to be particularly affected by deduplication: for example, Swedish loses approximately 16% and 18% of its documents to exact-match and MinHash, respectively. We interpret such findings to indicate that sheer data volume can be deceiving, and stress that NLP practitioners should be aware of such issues when utilizing Wikipedia as a resource.
4.2 Comparison to Depth+
In its official statistics, Wikimedia provides a Depth metric (alternatively, editing depth), designed to serve as a proxy measure for a given Wikipedia’s ‘collaborative quality’. [^8] https://meta.wikimedia.org/wiki/Wikipedia_article_depth However, Alshahrani et al. (2023) caution against utilizing it for such purposes. Specifically, they discuss the metric’s sensitivity to bot content (articles and edits) and excessive editing (‘edit wars’), as well as its inability to consolidate broader user activity and account for lower-resourced Wikipedias. As such, they propose a modified metric, Depth+:
$\mathrm{Depth}^{+}=\mathrm{Editors}\cdot\frac{\mathrm{Edits}}{\mathrm{Total}}\cdot\frac{\mathrm{Articles}}{\mathrm{Non-Articles}}$ (2)
Alshahrani et al. (2023) show that Depth+ is able to adequately account for bot-generated content, correctly de-ranking Wikipedias such as Vietnamese, Serbo-Croatian, and Cebuano. Conversely, they demonstrate that Depth+ likewise includes several low-resource Wikipedias, such as Cree, Tigre, and Bangla towards the top of the ranking, highlighting their users’ supposedly high degree of collaboration.
We find a significant positive correlation between our primary filtering procedure (measured as the percentage of documents retained) with Depth+ ($\rho=0.25,p<0.001$). Thus, there are similar notions of quality captured by both metrics. However, approximately half of all Wikipedias yield Depth+ values of less than 0.3, a majority very close to 0. The rest have a significantly higher Depth+ score (excluding English, German is highest with 36.04) resulting in many outliers that skew the distribution. We surmise that this is because the number of edits is disparagingly different across Wikipedia, effectively privileging Wikipedias where active users comprise a larger percentage of total users — a proxy for ‘collaborative quality’. However, this does not shed much insight on the quality of individual articles, written by one-time, potentially domain-expert authors.
4.3 Bot generation
We have observed that our primary filters remove a sizeable volume of articles from Wikipedias known to contain proliferate bot-generated content. To better understand if we are indeed filtering out such content, we refer to Alshahrani et al. (2023)’s estimations of bot-generated articles across Wikipedia. Indeed, we find that there is a strong negative correlation ($\rho=-0.56,p<0.001$) between the percentage of bot articles in a given Wikipedia and the percentage of articles remaining after primary filtering. We consider this to be an encouraging finding, provided that our primary filters are simple, pre-defined processes that state obvious assumptions about our ideal data distribution. Additionally, we do not rely on specialized bot-detection models trained on other distributions.
Figure 4: Percentage of total docs removed by MinHash and ratio of bot articles to total articles per Wikipedia.
Going further, we hypothesize that MinHash deduplication plays a large role in removing bot articles, given its ability fuzzy-match strings across documents. Here, we observe that the percentage of documents filtered out by MinHash yields an even higher correlation with bot article ratio ($\rho=0.63,p<0.001$) (Figure 4), indicating that this process is an adequate proxy for bot article removal in Wikipedia. Unsurprisingly, the Wikipedias with the highest percentage of bot content according to Alshahrani et al. (2023) — Cebuano (99%), Waray (90%), Swedish (68%) — tend to have much (if not most) of their content removed by MinHash.
4.4 Heuristics as Quality Measures
So far, we have largely focused on the effects of primary filters and their association with other recognized quality metrics. Here, explore how heuristic filters might likewise aid in our understanding of ‘quality’ as it is distributed across Wikipedia. To do so, we plot the distributions of each of the three heuristic classes defined in Table 1 across 8 Wikipedias - 2 from each tier in our quality taxonomy (Figure 5).
Immediately, we can observe that each heuristic class seems to possess its own distinct distributional type. For example, the absolute metrics tend to resemble a log-normal distribution, typically featuring a very short head and an exceedingly long tail. Intuitively, this makes sense, given that very short articles (e.g. less than ten characters) are likely to be rare, and that longer documents (but not too long) will be encountered with increasing frequency. Here, we can observe that both Hausa and Igbo (Tier 1) follow the expected pattern, featuring a short head and a gradually diminishing tail of long documents. Conversely, Waray and Minangkabau exhibit sharp, peaking heads with virtually no tail to follow them, indicating that these Wikipedias contain predominantly very short articles.
In contrast to the absolute metrics, the relative value distributions tend to be left-skewed, with the head comprising values higher than 0.5999Note that we plot combined and normalized metrics in the figure; each individual metric is bounded on $(0,1)$.. This is likewise expected, as low-quality articles for these metrics are likely to yield values closer to 0, indicating repetitive language use — for example, an article returning $\frac{3}{30}=0.1$ would feature one of three distinct tokens thirty times. Interestingly, this phenomenon is generally observed for all 8 languages, with the of exception of Amharic, which features a narrow, short mode around 0.5 — possibly indicating a cluster of low quality data.
Lastly, the entropy distributions tend to be somewhat normal in shape, which is indeed the desired outcome for this metric, signifying that documents follow predictable n-gram patterns. Generally, these metrics are somewhat related to the relative class, in that lower values will tend to correspond to highly predictable data. However, while higher-valued relative metrics are typically desirable, coinciding with varied lexical usage, higher-valued entropy values correspond instead to completely unpredictable — and potentially undecipherable — text. In the latter case, this could manifest by HTML or javascript leakage or poor page formatting. Interestingly, three of the four lower-tiered languages (Yoruba, Waray, Minangkabau) appear to exhibit a higher concentration of lower-entropy documents, indicating that they include more predictable, potentially bot-generated text. Conversely, Hausa and Igbo are left-skewed, implying that their articles feature more creative language use.
Figure 5: Distribution of select Wikipedias from Tier 1 (blue), Tier 2 (orange), Tier 3 (green) and Tier 4 (red) on each class of heuristic metrics.
5 Downstream Evaluation
Having demonstrated that our filtering procedures align with various qualitative notions of ‘quality’, we now turn to a quantitative evaluation focused on downstream NLP tasks. Specifically, we evaluate the effect of data pruning in three scenarios: monolingual pretraining, language adaptation, and multilingual pretraining (see Figure 6).
Figure 6: The experimental pipeline
5.1 Monolingual Pretraining
To isolate the impact of data quality on downstream performance, we pretrain monolingual masked language models on 50 Wikipedias from the 4 tiers described in Section 4, selected based on the availability of downstream evaluation tasks. We employ three different versions of these Wikipedias as pretraining data: (1) the entire Wikipedia of the language without any filtering (raw_wiki), (2) filtered Wikipedia using the primary filters (+primary), and (3) further pruned using the heuristic filters (+heuristic), as described in §3. In addition to this, we include a control experiment (random) wherein we sample $n$ random articles from the Wikipedia, where $n$ is equal to the number of articles filtered in the primary filtering step. The full experimental pipeline is illustrated in Figure 6. All models are subsequently fine-tuned on MasakhaNER 2.0 Adelani et al. (2022), MasakhaPOS Dione et al. (2023), MasakhaNEWS Adelani et al. (2023), AfriSenti-Twitter Muhammad et al. (2023) and SIB200 Adelani et al. (2024), based on availability.
We employ small monolingual models so as to minimize the risk of confounding factors such as unbalanced multilingual data or tokenisation inequalities, while gaining training efficiency (Warstadt et al., 2023). We use the DeBERTa architecture He et al. (2021), largely due to its competitive performance in the low-data regime (Samuel et al., 2023). Details about our model and tokeniser hyperparameters, and training regime can be found in Appendix A.
Figure 7: Results for monolingual pretraining across 4 tasks for +primary and +heuristic. The y-axis plots the respective performance difference for each filtering method, with the gray dotted line corresponding to no difference. The x-axis plots the fraction of data removed (of articles $\cdot$ % of characters), to visualise the interaction between data filtering and performance. Points are coloured to denote tier categories - blue (Tier 1), orange (Tier 2), green (Tier 3), red (Tier 4).
Table 2 displays the average scores for all languages across filter methods and tasks. Here, we can observe that, while +primary generally yields similar performance as raw_wiki (MasakhaNER and AfriSenti-Twitter), it occasionally leads to better performance than no filtering at all (MasakhaNews). With the exception of AfriSenti-Twitter, both +primary and +heuristic perform better than random, indicating that the quality filters capture a notion of quality that is beneficial to model performance. In general, primary filtering appears to be more effective on average than heuristic filtering.
Task raw_wiki random +primary +heuristic
NER 0.81 0.77 0.81 0.80
News 0.80 0.76 0.82 0.80
SentiA 0.59 0.59 0.59 0.59
SIB200 0.63 0.63 0.62 0.50
Table 2: Average scores for all languages across filtering methods and tasks. Individual results are Appendix A
We can note from Figure 7 that for MasakhaNER and MasakhaNEWS, the removal of a large amount of data does not lead to a performance decline — in fact, yo actually enjoys a performance increase, despite 65% of the Wikipedia being removed. Interestingly, most languages that see a slight performance decrease happen to be allocated to Tier 1 (blue), and thus lose a smaller fraction of the data. We observe that +heuristic elicits a similar trend for MasakhaNEWS, but likewise note that the general performance decline for MasakhaNER, with respect to raw_wiki.
Interestingly, we find that all filtering strategies — including random— yield identical performance on AfriSenti-Twitter. We surmise that this could be due to a severe domain mismatch between the formal, encyclopaedic nature of Wikipedia and the social media-oriented nature of this particular dataset. As such, our assumptions about quality pretraining data do not directly apply here. We do note that ary is an exception, where the raw_wiki and random (see Table 10) distributions outperform quality filtering to a large degree. This can be explained by the fact that the ary subset in AfriSenti-Twitter contains script switching between Arabic and Latin. In our training setup, only raw_wiki and random models are able to tokenise Latin characters due to the script filtering. This connects to the discussion in §2.3 about what data quality actually means (more in §6)
Figure 8: Distribution of raw_wiki (blue), +primary (orange), +heuristic (green) and random (red) scores with monolingual pretraining across 50 languages in SIB200.
With SIB200 we are able to evaluate 50 languages, evenly spread across tiers. Thus, we can adequately compare the effectiveness of data quality filtering with respect to the existing quality of the Wikipedia. A KDE plot of the distribution of scores for each filtering method separated by tier can be found in Figure 8. Here, we observe that the distribution of the random scores is, on average, lower than the other three methods — with a more pronounced difference seen in Tiers 2, 3 and 4 than Tier 1. The underlying distribution of the Wikipedia, thus, has a meaningful effect on the impact of the quality filtering — as we also observed with MasakhaNEWS and MasakhaNER. We consider this to be a notable finding, since primary and heuristic filtering remove a larger proportion of articles for lower-ranked Wikipedias than those in Tier 1. This implies that the articles removed are of poor quality and are not beneficial to the training process. As a testament to this, we point to ceb in Figure 7, which has a 12% performance gain with +primary. We attribute this primarily to the deduplication step, which removes most bot generated articles.
5.2 Language adaptation
While the small-scale models of the previous section grant us full control of crucial variables relevant to our study, we are likewise interested how methods fare in a more common scenario. To this end, we turn to language adaptation: a common practice in the realm of low-resource and cross-lingual NLP concerned with adapting existing mono or multilingual models to languages unseen during pretraining Pfeiffer et al. (2020, 2022). Here, we follow the experimental protocol of Alabi et al. (2022) (see Appendix A) in adapting XLM-R base (Conneau et al., 2020) to unseen languages from our experimental cohort. We use the original XLM-R tokeniser and further pretrain XLM-R on individual Wikipedias across the raw_wiki, +primary, and +heuristic settings. Following this, we fine-tune each model on SIB200 and Belebele Bandarkar et al. (2024) — a cross-lingual reading comprehension bencmark with coverage over 122 languages. [^10] We fine-tune the models on the English fine-tuning data recommended in the Belebele paper, and evaluate it on the target languages in a zero shot manner. This setup is not possible with the monolingual models due to tokeniser restrictions.
Figure 9: Performance difference for XLMR+primary and XLMR+heuristic relative to XLMR+raw_wiki on SIB200 (with in-language fine-tuning) and Belebele (zero-shot).
Figure 10: Performance difference for +primary and +heuristic with multilingual pretraining relative to raw_wiki.
In general, we find that results for SIB200 in Figure 9 echo those observed for the monolingual models in Figure 7. In other words - while all languages (except mni) benefit from language adaptation (see Table 12), the effect of quality filtering is dependent upon the underlying data distribution. Notably, Tier 2, Tier 3 and Tier 4 benefit more from primary filtering, with ace achieving an 8% performance boost. Tier 1 Wikipedias, on the other hand, do not see much performance improvement from filtering on average. We make a special note of mni, which suffers a staggering 16% performance drop after language adaptation (see Table 12), raising many questions about the quality of that particular Wikipedia. This gap is notably reduced with primary filtering, however.
We observe the greatest benefit of quality filtering — especially primary filtering — in the Belebele evaluation, depicted in Figure 9. With the exception of bm, all languages experience a substantial performance boost, amounting to 5% on average. Previous work has shown that quality of pretraining data is more relevant to performance in a zero-shot setting than when fine-tuning the model on the target language (Tatariya et al., 2023). Thus, the higher quality +primary data can be expected to provide a benefit here. Beyond this, we observe that +heuristic filtering does not provide an advantage over primary filtering in either task. However, model performance in this setting is still very close to raw_wiki, despite the large-scale article loss of various Wikipedias.
5.3 Multilingual Pretraining
In our last downstream experiment, we aim to evaluate the effect of data quality filtering in a multilingual pretraining scenario. To do so, we train three multilingual masked language models on the top 100 Wikipedias in terms of article count, mimicking the setup of multilingual BERT Devlin et al. (2019). As with the monolingual and language adaptation experiments, we partition our data according to the raw_wiki, +primary and +heuristic settings. Across all models, all hyperparameters (see Appendix A) are kept constant, and each model is trained for the same amount of steps, allowing for comparability across settings. After training, we fine-tune these models on English task-specific data and evaluate zero-shot on Belebele and XNLI (Conneau et al., 2018).
Ultimately, we observe similar trends as before on Belebele (Figure 10), with the lower tiers showing more performance gains than Tier 1 with +primary. However, we do not observe the same degree of performance increase as depicted in Figure 9. For example, while ht enjoyed a 9% performance gain with XLMR in the +primary setting, it remains on par with raw_wiki after multilingual pretraining. On the other hand, other languages such as yo do not demonstrate any performance difference across settings. We interpret this to indicate that factors other than quality data can affect downstream performance for multilingual models, such as the language choice and composition. Going further, we note that +heuristic results also follow trends from earlier experiments, with most languages clustered around the dotted line — i.e. similar performance to raw_wiki. Notably, we find that war and ceb appear to benefit from the +heuristic step, despite retaining only on 10% and 18% of their data, respectively. With XNLI, we see that performance generally degrades after both quality filtering settings. However, it is important to note that most languages in XNLI have Wikipedias categorized under Tier 1. Notably, the performance for vi remains close to raw_wiki for both filtering methods, continuing to follow the overall trend.
6 Discussion and Recommendations
Effect of Quality Filtering
The quality filters tested in our experiments outperform random filtering, thus establishing that the primary and heuristic filters capture a notion of quality relevant to downstream performance. Performance after data quality filtering generally remains on par or better than no filtering at all. We look for a correlation between data pruning and accuracy with a Spearman’s correlation test between the amount of data removed in the primary filtering step, and the accuracy difference between raw_wiki and +primary. There is no significant correlation, which means that data quality pruning may not have a direct significant impact on performance. Findings from van Noord et al. (2024) and Artetxe et al. (2022) also echo this result. A substantial amount of bad data can be pruned without seeing any performance loss. In fact, Wikipedias in Tier 2, Tier 3 and Tier 4 benefit more from primary filtering than those in Tier 1. Thus, we recommend data quality filtering as a means for resource efficient training, in addition to the benefit of better performance for many languages.
Heuristic filtering, on the other hand, yields mixed results. This cannot be entirely attributed to aggressive data filtering, as Tier 1 Wikipedias have more performance decline than those in lower tiers. We speculate that this discrepancy arises because heuristic metrics are computed at the article level, whereas model training occurs at the chunk level. Likewise, the heuristic filters that we deploy can be considered rather generic in nature, accounting only for distributionally-conditioned word and character-level statistics. It is likely that, for language-specific applications, one may include specialized heuristics that capture more information about the data domain — e.g. specific keywords or phrases — and choose specific thresholds for these purposes. Furthermore, +primary may already possess a well-balanced data distribution, making further filtering detrimental. Regardless, we would like to note that heuristic filtering still has its merits for many languages, thus warranting further analysis in future work.
Between monolingual and multilingual pretraining, the advantages of quality filtering are seen most in the monolingual setting, especially in the zero-shot scenario. Trends noted above remain observable in the multilingual models but less pronounced, likely due to cross-lingual interactions influencing the outcomes. We recommend always using data quality filters when fine-tuning data is not available, or when adapting existing models to a low-resource language, especially if the Wikipedia is in Tier 2, Tier 3 and Tier 4.
Quality versus Size
Figures 7, 9 and 10 demonstrate that primary and heuristic filtering steps can potentially prune up to 60-80% of the data in a bad Wikipedia. This is also observed for lower resource languages with small Wikipedias (see yo, nso, lg, pag, to name a few) where the assumption might be that there is already very little data available, rendering any further filtering to be detrimental. We demonstrate that even for such small Wikipedias, removal of bad data is desirable, even beneficial to performance on downstream tasks. Moreover, when compared with the XLM-R model adapted to the same data (for e.g., monolingual raw_wiki versus XLM-R+raw_wiki), 23 monolingual models either match or outperform XLM-R on SIB200 — all with a fraction of the data size and capacity. Moreover, our results for AfriSenti-Twitter are competitive with those reported by Muhammad et al. (2023) on larger multilingual models. This indicates that a smaller concentration of monolingual data trained with fewer parameters can yield better performance than a large multilingual corpus that includes the same data. Therefore, low-resource languages benefit from smaller, more efficient models trained on a fraction of the compute with quality filtering being paramount to achieving good performance and getting the most out of limited data.
Quality versus Language
Although the results discussed in this study largely remain consistent in aggregate, zooming in on individual languages reveals the large benefits of the filtering process for some languages. For example, om and ti benefit from a substantial performance boost on MasakhaNEWS after primary filtering. Although +heuristic generally does not prove to be as competitive as +primary, certain languages — such as war and jv with multilingual pretraining — still benefit considerably from the process. As such, we emphasize that there there exists no one-size-fits-all data filtering approach. Instead, we advocate for practitioners to move away from seeking general definitions of data quality and focus rather on the examination of filtering strategies for individual Wikipedias. Concretely, data quality decisions with regard to user-generated multilingual corpora such as Wikipedia should be made on a per-language basis, despite the theoretical domain to which they belong. Furthermore, this approach can benefit from leveraging typological features and databases to create more language-specific data filtering pipelines, which is a promising direction for future work.
Quality and Downstream Tasks
Aside from a pretraining resource, Wikipedia is an equally important primary source for multilingual, task-specific dataset creation. While we have focused largely on the former, we highlight some implications of data quality filtering for the latter.
Lent et al. (2024) examined the quality of Creole Wikipedias, observing that many articles contain short and templatic entries as a result of machine generation. In this work, we have engaged in discussions regarding the benefit of removing such articles in a pretraining setting. However, Lent et al. (2024) note that the templatic nature of such Wikipedias, in fact, happen to facilitate their annotation process in creating a relation-classification dataset. In such scenarios, our filtering pipeline can be employed to describe and identify useful articles. For example, in this case Lent et al. (2024)’s experiments, heuristic filtering can be employed to identify articles that are very short, with low entropy, and thus are more likely to be templatic and generated from a knowledge base.
In contrast to this, Kreutzer et al. (2022a) discuss the pervasive impact of low data quality in downstream evaluation with the example of machine translation. This is particularly salient in the low-resource regime, where Wikipedia is often the only source of data — potentially leading to flawed evaluation datasets and misleading conclusions. Lignos et al. (2023) highlight quality issues with the WikiAnn dataset Pan et al. (2017), reporting much noise present in the data, along with many examples that either are very dense in mentions or consist entirely of single entity mentions, particularly for African languages. This could likely stem from the broader quality issues endemic to Wikipedia.
7 Conclusion
In this study, we challenged the assumption of Wikipedia as a high quality resource in multilingual NLP, and provided a guide to utilizing it as a pretraining resource in NLP. We show that much of Wikipedia contains numerous quality issues such as duplicate articles, bot-generated content, and foreign scripts, among other issues, which can be filtered out as a means to achieve good performance with less compute. For select languages, this process can even lead to performance improvement. Overall, we advocate towards a shift in perspective from seeking general definitions of data quality to approaching quality issues from a domain or language specific lens.
8 Acknowledgements
For KT, AK, WP and MDL the computational resources and services used were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government - department EWI. AK was partially funded by a Google Research Award. WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. EP, HL and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF210454). We would also like to thank Isaac Caswell, Colin Cherry, and Markus Freitag at Google Translate Research for their insightful reviews on the initial draft.
Appendix A Training Hyperparameters and Finetuning Results
Hyperparameters for the pretraining and fine-tuning of monolingual models are in Table 3. We use the SentencePiece package Kudo and Richardson (2018) to train monolingual unigram tokenisers for each language. We train two tokenisers per language: one on the raw_wiki with no filtering, which is used to train the raw_wiki models; the second one on the data remaining after script filtering and deduplication, which we use to train all other models. We use an approximation with Zipf’s law Zipf (1949) on each Wikipedia to estimate the vocabulary size for each model, mentioned in Table 6. Hyperparameters for XLM-R language adaptation are in Table 4. For both cases, we checkpoint and use the best model on a left-out validation set to fine-tune after pretraining. For multilingual pretraining, we train 2 WordPiece tokenizers on raw_wiki and +primary (used to train both +primary and +heuristic models) with a vocabulary size of 110k. Languages are listed in Table 7 and hyperparameters in Table 5.
Tiny DeBERTa Hyperparameters
Pretraining
Model Type deberta
Hidden Size 312
Intermediate Size 1200
Attention Heads 12
Hidden Layers 4
Hidden Act gelu
Max Length 128
Task MLM
Mask Probability 0.4
Padding Strategy longest
Grad. Accumulation Steps 4
Batch size 8 to 32
Learning Rate 1.00e-3
Epochs 100
Finetuning
Max Length 128
Epochs 20
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Table 3: Hyperparameters for monolingual pretraining and fine-tuning. While we used 100 epochs as a default, for Wikipedias that are too large (nl, ro, id, hu, uk, vi, ceb, war, tt, azb, min) we pretrain only for 2 epochs. Similarly, while we used a batch size of 8 for all models, for the larger Wikipedias we use a batch size of 32. All models are trained on 1 GPU.
XLMR Hyperparameters
Language Adaptation
Model Type xlm-roberta
Task MLM
Max Length 512
Mask Probability 0.15
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding strategy Longest
Grad. Accumulation Steps 1
Finetuning - SIB200
Max Length 512
Epochs 20
Batch Size 32
Learning Rate 5.00e-5
Padding Strategy max_length
Finetuning - Belebele
Max Length 128
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Table 4: Hyperparameters for XLM-R language adaptation (based on Alabi et al. (2022)) and fine-tuning.
Multilingual Pretraining
Pretraining
Sampling Temperature 1.43
Model Type bert
Task MLM
Max Length 128
Mask Probability 0.12
Steps 200,000
Batch Size (per device) 256
Learning Rate 1.00e-3
Padding strategy Longest
Grad. Accumulation Steps 4
Finetuning - Belebele
Max Length 128
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Finetuning - MNLI
Max Length 128
Epochs 3
Batch Size 32
Learning Rate 2.00e-5
Padding Strategy max_length
Table 5: Hyperparameters for multilingual pretraining and fine-tuning, based on mBERT and Izsak et al. (2021). All three models are trained on 4 H100 GPUs.
Estimated Vocabulary Size
Lang Vocab Wiki-ID Vocab
ace 3505 mr 14708
am 5234 nl 60011
ary 4502 nso 2801
awa 2190 ny 2472
azb 13996 om 3436
ban 6624 pag 2170
bm 1485 pcm 2595
ceb 73124 rn 1446
ckb 11451 ro 36329
ee 1770 rw 5725
fon 1426 shn 5784
ha 13126 sn 5227
ht 10836 so 6117
hu 46205 sw 12529
id 41725 te 23175
ig 11609 ti 1150
kk 22496 tn 3783
ks 2248 ts 1817
lg 4528 tt 26047
lij 5462 tw 4745
ln 2509 uk 62626
mai 4793 vi 43806
mi 3451 war 25803
min 14783 wo 3384
mni 3351 yo 6128
zu 4439
Table 6: Estimated vocabulary size used to train monolingual tokenisers for each Wikipedia in our experimental pipeline.
Languages Included in Multilingual Pretraining
Language Wiki-ID Language Wiki-ID
Afrikaans af Korean ko
Albanian sq Latin la
Arabic ar Latvian lv
Aragonese an Lithuanian lt
Armenian hy Lombard lmo
Asturian ast Low German nds-nl
Azerbaijani az Luxembourgish lb
Bashkir ba Macedonian mk
Basque eu Malagasy mg
Bavarian bar Malay ms
Belarusian be Malayalam ml
Bangla bn Marathi mr
Bishnupriya bpy Minangkabau min
Bosnian bs Nepali ne
Breton br Newari new
Bulgarian bg Norwegian Nynorsk nn
Burmese my Occitan oc
Catalan ca Persian fa
Chechen ce Piedmontese pms
Chinese zh Polish pl
Cantonese zh-yue Portuguese pt
Cebuano ceb Punjabi pa
Chuvash cv Romanian ro
Croatian hr Rusyn rue
Czech cs Scots sco
Danish da Serbo-Croatian sh
Dutch nl Sicilian scn
English en Slovak sk
Estonian et Slovenian sl
Finnish fi South Azerbaijani azb
French fr Spanish es
Galician gl Sundanese su
Georgian ka Swahili sw
German de Swedish sv
Greek el Filipino tl
Gujarati gu Tajik tg
Haitian Creole ht Tamil ta
Hebrew he Tatar tt
Hindi hi Turkish tr
Hungarian hu Ukrainian uk
Icelandic is Uzbek uz
Ido io Vietnamese vi
Indonesian id Volapük vo
Irish ga Waray war
Italian it Welsh cy
Japanese ja Western Frisian fy
Javanese jv Western Panjabi pnb
Kannada kn Yoruba yo
Kazakh kk Thai th
Kyrgyz ky Mongolian mn
Table 7: Wikipedias included in the pretraining of the multilingual models.
MasakhaNER 2
Lang raw_wiki +primary +heuristic random
pcm 0.77±0.00 0.76±0.00 0.76±0.00 0.74±0.00
ha 0.82±0.00 0.81±0.00 0.81±0.00 0.75±0.01
ig 0.85±0.00 0.85±0.00 0.84±0.00 0.81±0.01
sn 0.92±0.00 0.92±0.00 0.92±0.00 0.90±0.00
rw 0.76±0.05 0.79±0.01 0.77±0.05 0.74±0.01
tw 0.76±0.00 0.76±0.01 0.74±0.00 0.71±0.00
wo 0.75±0.01 0.73±0.01 0.72±0.01 0.73±0.01
ny 0.87±0.00 0.87±0.00 0.86±0.00 0.85±0.00
ee 0.81±0.00 0.81±0.01 0.79±0.00 0.78±0.01
bm 0.74±0.01 0.74±0.01 0.72±0.01 0.73±0.00
fon 0.74±0.01 0.74±0.00 0.76±0.01 0.71±0.01
sw 0.90±0.00 0.90±0.00 0.90±0.00 0.88±0.00
zu 0.83±0.01 0.82±0.01 0.81±0.01 0.77±0.01
lg 0.84±0.00 0.85±0.00 0.83±0.00 0.80±0.01
tn 0.80±0.01 0.80±0.00 0.79±0.01 0.63±0.01
yo 0.84±0.00 0.84±0.00 0.83±0.00 0.80±0.00
Avg 0.81 0.81 0.8 0.77
Table 8: Fine-tuning results on MasakhaNER 2. Results are averaged over 5 runs, and languages are grouped by tier.
MasakhaNEWS
wiki raw_wiki +primary +heuristic random
pcm 0.90±0.04 0.91±0.00 0.87±0.04 0.88±0.02
ha 0.88±0.01 0.86±0.01 0.86±0.01 0.84±0.01
ig 0.85±0.01 0.85±0.01 0.84±0.01 0.81±0.01
sn 0.89±0.01 0.90±0.01 0.89±0.01 0.87±0.01
so 0.71±0.01 0.71±0.01 0.70±0.01 0.59±0.03
ln 0.78±0.01 0.76±0.01 0.76±0.01 0.74±0.02
om 0.67±0.30 0.82±0.01 0.81±0.30 0.71±0.04
ti 0.60±0.04 0.65±0.02 0.62±0.04 0.56±0.02
sw 0.81±0.01 0.82±0.00 0.82±0.01 0.75±0.01
am 0.84±0.04 0.87±0.01 0.85±0.04 0.83±0.01
lg 0.86±0.01 0.86±0.02 0.84±0.01 0.84±0.02
rn 0.75±0.01 0.76±0.02 0.71±0.01 0.71±0.01
yo 0.84±0.03 0.86±0.01 0.86±0.03 0.82±0.01
Avg 0.8 0.82 0.8 0.76
Table 9: Fine-tuning results on MasakhaNEWS. Results are averaged over 5 runs, and languages are grouped by tier.
AfriSenti-Twitter
wiki raw_wiki +primary +heuristic random
pcm 0.61±0.01 0.61±0.01 0.61±0.01 0.60±0.01
ha 0.72±0.01 0.72±0.01 0.73±0.01 0.72±0.01
ig 0.74±0.00 0.75±0.00 0.75±0.00 0.74±0.01
rw 0.60±0.04 0.61±0.01 0.59±0.04 0.60±0.01
tw 0.62±0.01 0.64±0.01 0.64±0.01 0.63±0.01
ts 0.48±0.04 0.49±0.03 0.50±0.04 0.50±0.01
sw 0.55±0.03 0.59±0.02 0.56±0.03 0.53±0.04
am 0.40±0.07 0.41±0.04 0.46±0.07 0.45±0.10
ary 0.46±0.01 0.40±0.01 0.40±0.01 0.44±0.01
yo 0.68±0.01 0.68±0.01 0.68±0.01 0.64±0.01
Avg 0.59 0.59 0.59 0.59
Table 10: Fine-tuning results on AfriSenti-Twitter. Results are averaged over 5 runs, and languages are grouped by tier.
SIB200
wiki raw_wiki +primary +heuristic random
Tier 1
ti 0.38±0.05 0.38±0.01 0.38±0.05 0.32±0.01
ha 0.77±0.02 0.78±0.02 0.75±0.02 0.62±0.03
ts 0.61±0.03 0.60±0.01 0.56±0.03 0.53±0.02
ig 0.76±0.01 0.77±0.02 0.74±0.01 0.64±0.02
sn 0.66±0.03 0.62±0.03 0.61±0.03 0.43±0.05
so 0.68±0.01 0.73±0.02 0.70±0.01 0.56±0.06
rw 0.77±0.01 0.77±0.01 0.73±0.01 0.65±0.03
tw 0.65±0.10 0.65±0.02 0.67±0.10 0.63±0.03
ln 0.59±0.02 0.57±0.02 0.56±0.02 0.52±0.04
wo 0.50±0.02 0.52±0.02 0.45±0.02 0.45±0.04
om 0.59±0.01 0.62±0.01 0.56±0.01 0.34±0.05
ny 0.64±0.02 0.60±0.01 0.64±0.02 0.60±0.01
ee 0.59±0.03 0.61±0.03 0.57±0.03 0.57±0.01
bm 0.45±0.02 0.48±0.02 0.47±0.02 0.47±0.03
fon 0.48±0.03 0.45±0.04 0.48±0.03 0.44±0.01
Avg 0.61 0.61 0.59 0.51
Tier 2
ro 0.82±0.03 0.80±0.02 0.81±0.03 0.66±0.01
id 0.83±0.01 0.85±0.01 0.85±0.01 0.75±0.04
hu 0.82±0.01 0.86±0.01 0.82±0.01 0.64±0.05
nl 0.82±0.03 0.80±0.03 0.82±0.03 0.66±0.03
uk 0.85±0.01 0.81±0.02 0.84±0.01 0.43±0.08
awa 0.54±0.03 0.52±0.03 0.54±0.03 0.47±0.02
ks 0.42±0.02 0.45±0.02 0.42±0.02 0.41±0.05
ckb 0.80±0.01 0.82±0.01 0.82±0.01 0.54±0.04
rn 0.51±0.03 0.50±0.02 0.51±0.03 0.45±0.03
sw 0.79±0.02 0.81±0.02 0.83±0.02 0.58±0.05
am 0.66±0.01 0.65±0.02 0.63±0.01 0.41±0.02
zu 0.73±0.01 0.70±0.02 0.69±0.01 0.61±0.02
tn 0.64±0.02 0.61±0.01 0.64±0.02 0.37±0.05
ary 0.59±0.00 0.60±0.01 0.61±0.00 0.46±0.02
lg 0.55±0.02 0.55±0.02 0.56±0.02 0.51±0.02
Avg 0.69 0.69 0.69 0.53
Tier 3
lij 0.61±0.01 0.59±0.02 0.54±0.01 0.45±0.02
ht 0.70±0.04 0.69±0.01 0.66±0.04 0.36±0.03
te 0.80±0.02 0.79±0.02 0.81±0.02 0.52±0.03
mr 0.74±0.04 0.78±0.02 0.76±0.04 0.50±0.04
kk 0.81±0.02 0.83±0.02 0.80±0.02 0.74±0.02
yo 0.70±0.02 0.69±0.02 0.70±0.02 0.52±0.07
ace 0.27±0.02 0.15±0.03 0.17±0.02 0.23±0.01
shn 0.74±0.02 0.72±0.02 0.75±0.02 0.62±0.04
mai 0.67±0.01 0.68±0.02 0.68±0.01 0.55±0.02
ban 0.65±0.01 0.66±0.02 0.68±0.01 0.49±0.03
vi 0.82±0.01 0.83±0.00 0.81±0.01 0.74±0.02
nso 0.50±0.02 0.53±0.01 0.46±0.02 0.46±0.04
Avg 0.67 0.66 0.65 0.52
Tier 4
tt 0.79±0.01 0.75±0.01 0.76±0.01 0.48±0.03
ceb 0.39±0.03 0.51±0.03 0.50±0.03 0.28±0.04
war 0.48±0.04 0.49±0.04 0.50±0.04 0.46±0.02
pag 0.45±0.03 0.48±0.01 0.46±0.03 0.35±0.02
mi 0.55±0.02 0.52±0.02 0.52±0.02 0.48±0.03
mni 0.27±0.04 0.21±0.01 0.21±0.04 0.31±0.00
azb 0.53±0.05 0.53±0.02 0.51±0.05 0.42±0.03
min 0.65±0.03 0.60±0.03 0.53±0.03 0.41±0.02
Avg 0.51 0.51 0.50 0.40
Table 11: Fine-tuning results on SIB200 for tiny DeBERTa models. Results are averaged over 5 runs.
XLMR Continued Pretraining (SIB200)
wiki baseline raw_wiki +primary +heuristic
ti 0.59±0.04 0.53±0.21 0.56±0.02 0.60±0.21
fon 0.43±0.20 0.60±0.04 0.56±0.02 0.57±0.04
ig 0.52±0.20 0.79±0.02 0.79±0.03 0.75±0.02
ee 0.64±0.04 0.66±0.03 0.64±0.03 0.64±0.03
ny 0.64±0.01 0.68±0.03 0.65±0.02 0.68±0.03
wo 0.60±0.03 0.65±0.01 0.64±0.01 0.65±0.01
ln 0.66±0.01 0.70±0.01 0.69±0.01 0.68±0.01
tw 0.66±0.02 0.72±0.02 0.70±0.03 0.70±0.02
rw 0.56±0.03 0.69±0.02 0.67±0.02 0.63±0.02
sn 0.54±0.03 0.59±0.02 0.60±0.03 0.60±0.02
bm 0.43±0.22 0.60±0.05 0.55±0.03 0.54±0.05
ts 0.58±0.05 0.63±0.04 0.60±0.05 0.59±0.04
Avg 0.57 0.65 0.64 0.64
tn 0.56±0.02 0.62±0.03 0.60±0.04 0.55±0.03
ks 0.62±0.22 0.68±0.02 0.68±0.02 0.68±0.02
rn 0.54±0.04 0.55±0.06 0.58±0.02 0.57±0.06
lg 0.40±0.17 0.58±0.03 0.58±0.03 0.55±0.03
zu 0.58±0.01 0.64±0.03 0.68±0.02 0.65±0.03
awa 0.80±0.04 0.81±0.02 0.83±0.03 0.80±0.02
Avg 0.59 0.65 0.66 0.63
ht 0.63±0.03 0.73±0.03 0.73±0.02 0.72±0.03
nso 0.55±0.02 0.59±0.04 0.63±0.02 0.57±0.04
lij 0.77±0.02 0.81±0.03 0.78±0.03 0.80±0.03
ace 0.41±0.09 0.45±0.13 0.53±0.06 0.51±0.13
shn 0.33±0.16 0.61±0.03 0.60±0.04 0.62±0.03
mai 0.82±0.02 0.83±0.02 0.82±0.02 0.85±0.02
ban 0.79±0.01 0.82±0.02 0.83±0.01 0.83±0.02
yo 0.56±0.02 0.63±0.04 0.66±0.02 0.62±0.04
Avg 0.61 0.68 0.70 0.69
mi 0.59±0.12 0.62±0.02 0.63±0.03 0.59±0.02
mni 0.54±0.05 0.38±0.11 0.45±0.03 0.45±0.11
pag 0.77±0.02 0.79±0.02 0.77±0.05 0.78±0.02
Avg 0.59 0.66 0.66 0.65
Table 12: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on SIB200. Results are grouped by tier.
XLMR Continued Pretraining (Belebele)
wiki baseline raw_wiki +primary +heuristic
wo 0.27 0.26 0.29 0.26
ts 0.24 0.25 0.33 0.25
ny 0.26 0.25 0.27 0.25
rw 0.24 0.23 0.30 0.27
ig 0.26 0.25 0.30 0.22
ln 0.25 0.24 0.30 0.24
sn 0.25 0.23 0.33 0.22
bm 0.24 0.26 0.24 0.25
ti 0.24 0.27 0.29 0.24
lg 0.24 0.24 0.28 0.26
tn 0.25 0.24 0.33 0.25
zu 0.25 0.23 0.32 0.27
shn 0.25 0.26 0.27 0.24
nso 0.25 0.26 0.31 0.26
ht 0.24 0.24 0.33 0.25
yo 0.26 0.25 0.27 0.23
mi 0.26 0.27 0.29 0.24
Avg 0.25 0.25 0.30 0.25
Table 13: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on MC and tested on belebele. Languages are grouped by tier.
Multilingual Pretraining (XNLI)
wiki raw_wiki +primary +heuristic
el 0.39 0.39 0.39
bg 0.42 0.40 0.38
de 0.46 0.43 0.43
tr 0.43 0.41 0.41
th 0.38 0.37 0.38
fr 0.48 0.46 0.47
ru 0.40 0.38 0.37
es 0.50 0.49 0.50
zh 0.43 0.43 0.39
en 0.72 0.71 0.70
ar 0.39 0.38 0.38
sw 0.41 0.39 0.40
ur 0.38 0.39 0.37
hi 0.38 0.37 0.36
vi 0.41 0.41 0.41
Avg 0.44 0.43 0.42
Table 14: Multilingual models pretrained on top 100 Wikipedias with raw_wiki, +primary and +heuristic fine-tuned on MNLI and tested on XNLI. Languages are grouped by tier.
Multilingual Pretraining (Belebele)
wiki raw_wiki +primary +heuristic
pa 0.27 0.26 0.27
ta 0.24 0.23 0.25
af 0.24 0.23 0.25
th 0.22 0.25 0.24
fr 0.24 0.26 0.25
es 0.22 0.26 0.24
cs 0.26 0.24 0.25
pl 0.26 0.24 0.25
de 0.27 0.26 0.26
en 0.25 0.26 0.25
fi 0.26 0.25 0.26
pt 0.25 0.27 0.25
hr 0.25 0.27 0.27
el 0.24 0.23 0.23
da 0.26 0.25 0.24
it 0.25 0.26 0.24
bn 0.25 0.27 0.25
ml 0.27 0.24 0.25
is 0.25 0.27 0.27
bg 0.27 0.24 0.25
kn 0.26 0.25 0.25
et 0.27 0.26 0.25
ca 0.28 0.26 0.28
ko 0.27 0.20 0.25
tr 0.24 0.26 0.23
he 0.27 0.24 0.27
zh 0.25 0.24 0.24
sl 0.25 0.25 0.26
ja 0.27 0.24 0.23
lt 0.26 0.23 0.24
sk 0.23 0.24 0.25
ky 0.24 0.26 0.26
ka 0.25 0.25 0.26
mk 0.27 0.25 0.26
uk 0.24 0.26 0.23
hu 0.25 0.27 0.27
ro 0.25 0.27 0.24
id 0.26 0.27 0.27
hi 0.26 0.24 0.27
jv 0.23 0.24 0.29
hy 0.26 0.24 0.23
my 0.23 0.24 0.23
nl 0.26 0.23 0.24
sv 0.22 0.26 0.25
yo 0.23 0.26 0.25
vi 0.24 0.26 0.22
tg 0.24 0.24 0.22
su 0.24 0.28 0.26
gu 0.23 0.24 0.24
eu 0.26 0.25 0.25
ht 0.26 0.26 0.25
kk 0.27 0.26 0.27
mr 0.28 0.24 0.23
ceb 0.25 0.26 0.26
war 0.24 0.23 0.26
Avg 0.25 0.25 0.25
Table 15: Multilingual models pretrained on top 100 Wikipedias with raw_wiki, +primary and +heuristic fine-tuned on MC and tested on belebele. Languages are grouped by tier.
Appendix B Thresholding Details
We detail the procedure for creating thresholds for heuristic metrics as follows: for a given Wikipedia metric distribution (e.g., rps_doc_word_count), we determine a representative sample size based on the full dataset. Testing across various metrics and Wikipedias, we select a sample size of $n_{\mathrm{sample}}=0.05*n_{\mathrm{docs}}$, where $n_{\mathrm{docs}}$ is the total number of documents in the dataset. We then sort the full distribution in ascending order and take the first $n_{\mathrm{sample}}$ values as the value distribution, $D_{\mathrm{value}}$. This distribution reflects the lowest values in the distribution —essentially, the values we want to assess as potential outliers or benchmarks. Separately, we randomly sample $n_{\mathrm{sample}}$ values from the full distribution to create the sampled distribution, $D_{\mathrm{sample}}$. This random sample provides a baseline density that represents the overall data distribution rather than just extreme values. We estimate the kernel density (KDE) for both $D_{\mathrm{value}}$ and $D_{\mathrm{sample}}$, applying these KDEs over $n_{\mathrm{sample}}$ evenly spaced values in the range:
$\left[\min(D_{\mathrm{value}}),\,\max(D_{\mathrm{sample}})\right]$
The threshold is then identified as the point within this range where the difference between the two density estimates is maximized.
Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier Wiki-Id Tier
ab Tier 3 crh Tier 3 ha Tier 1 lez Tier 1 pa Tier 1 su Tier 3
ace Tier 3 cs Tier 1 hak Tier 3 lfn Tier 1 pag Tier 4 sv Tier 3
ady Tier 1 csb Tier 1 haw Tier 3 lg Tier 2 pam Tier 2 sw Tier 2
af Tier 1 cu Tier 1 he Tier 1 li Tier 1 pap Tier 1 szl Tier 3
als Tier 1 cv Tier 4 hi Tier 2 lij Tier 3 pcd Tier 1 ta Tier 1
alt Tier 2 cy Tier 2 hif Tier 2 lld Tier 4 pcm Tier 1 tay Tier 2
am Tier 2 da Tier 1 hr Tier 1 lmo Tier 3 pdc Tier 1 tcy Tier 1
ami Tier 2 dag Tier 3 hsb Tier 2 ln Tier 1 pfl Tier 3 te Tier 3
an Tier 1 de Tier 1 ht Tier 3 lo Tier 2 pi Tier 4 tet Tier 1
ang Tier 1 din Tier 1 hu Tier 2 lt Tier 1 pih Tier 1 tg Tier 3
ar Tier 2 diq Tier 3 hy Tier 2 ltg Tier 1 pl Tier 1 th Tier 1
arc Tier 1 dsb Tier 1 hyw Tier 1 lv Tier 1 pms Tier 3 ti Tier 1
arz Tier 3 dty Tier 2 ia Tier 3 mad Tier 1 pnb Tier 2 tk Tier 2
as Tier 1 dv Tier 1 id Tier 2 mai Tier 3 pnt Tier 2 tl Tier 1
ast Tier 2 dz Tier 1 ie Tier 3 map-bms Tier 3 ps Tier 1 tly Tier 3
atj Tier 2 ee Tier 1 ig Tier 1 mdf Tier 3 pt Tier 1 tn Tier 2
av Tier 2 el Tier 1 ik Tier 1 mg Tier 3 pwn Tier 1 to Tier 1
awa Tier 2 eml Tier 3 ilo Tier 2 mhr Tier 2 qu Tier 2 tpi Tier 2
ay Tier 1 eo Tier 1 inh Tier 1 mi Tier 4 rm Tier 2 tr Tier 1
az Tier 2 es Tier 1 io Tier 2 min Tier 4 rmy Tier 1 trv Tier 1
azb Tier 4 et Tier 1 is Tier 1 mk Tier 2 rn Tier 2 ts Tier 1
ba Tier 2 eu Tier 3 it Tier 1 ml Tier 1 ro Tier 2 tt Tier 4
ban Tier 3 ext Tier 1 iu Tier 2 mn Tier 1 roa-rup Tier 1 tum Tier 3
bar Tier 1 fa Tier 2 ja Tier 1 mni Tier 4 roa-tara Tier 2 tw Tier 1
bat-smg Tier 3 fat Tier 1 jam Tier 1 mnw Tier 1 ru Tier 1 ty Tier 3
bcl Tier 2 ff Tier 1 jbo Tier 1 mr Tier 3 rue Tier 2 tyv Tier 1
be Tier 2 fi Tier 1 jv Tier 2 mrj Tier 2 rw Tier 1 udm Tier 2
be-x-old Tier 1 fiu-vro Tier 1 ka Tier 2 ms Tier 3 sa Tier 2 ug Tier 2
bg Tier 1 fj Tier 1 kaa Tier 1 mt Tier 1 sah Tier 1 uk Tier 2
bh Tier 2 fo Tier 1 kab Tier 1 mwl Tier 1 sat Tier 1 ur Tier 2
bi Tier 1 fon Tier 1 kbd Tier 1 my Tier 2 sc Tier 1 uz Tier 2
bjn Tier 3 fr Tier 1 kbp Tier 1 myv Tier 2 scn Tier 2 ve Tier 3
blk Tier 1 frp Tier 1 kcg Tier 1 mzn Tier 3 sco Tier 1 vec Tier 3
bm Tier 1 frr Tier 1 kg Tier 2 nah Tier 2 sd Tier 1 vep Tier 1
bn Tier 1 fur Tier 1 ki Tier 2 nap Tier 3 se Tier 2 vi Tier 3
bo Tier 2 fy Tier 1 kk Tier 3 nds Tier 1 sg Tier 1 vls Tier 1
bpy Tier 3 ga Tier 1 kl Tier 1 ne Tier 1 sh Tier 3 vo Tier 2
br Tier 1 gag Tier 2 km Tier 1 new Tier 4 shi Tier 2 wa Tier 1
bs Tier 2 gan Tier 3 kn Tier 1 nia Tier 2 shn Tier 3 war Tier 4
bug Tier 4 gcr Tier 1 ko Tier 1 nl Tier 2 si Tier 1 wo Tier 1
bxr Tier 1 gd Tier 1 koi Tier 3 nn Tier 1 simple Tier 1 wuu Tier 1
ca Tier 1 gl Tier 1 krc Tier 2 no Tier 1 sk Tier 2 xal Tier 3
cbk-zam Tier 3 glk Tier 3 ks Tier 2 nov Tier 2 sl Tier 1 xh Tier 2
cdo Tier 3 gn Tier 1 ksh Tier 1 nrm Tier 2 sm Tier 1 xmf Tier 2
ce Tier 4 gom Tier 1 ku Tier 3 nso Tier 3 sn Tier 1 yi Tier 1
ceb Tier 4 gor Tier 3 kv Tier 2 nv Tier 3 so Tier 1 yo Tier 3
ch Tier 2 got Tier 2 kw Tier 1 ny Tier 1 sq Tier 2 za Tier 2
chr Tier 3 gu Tier 3 ky Tier 2 oc Tier 2 sr Tier 3 zea Tier 2
chy Tier 1 guc Tier 1 la Tier 2 olo Tier 1 srn Tier 3 zh Tier 1
ckb Tier 2 gur Tier 1 lad Tier 1 om Tier 1 ss Tier 1 zh-classical Tier 1
co Tier 1 guw Tier 1 lb Tier 1 or Tier 2 st Tier 2 zh-min-nan Tier 4
cr Tier 3 gv Tier 1 lbe Tier 2 os Tier 2 stq Tier 1 zh-yue Tier 2
zu Tier 2
Table 16: Quality categorisations for all non-English Wikipedias.